### v1

# Language-Integrated Recursive Queries

Abstract.
Performance-critical industrial applications, including large-scale program, network, and distributed system analyses, rely on fixed-point computations. The introduction of recursive common table expressions (CTEs) using the WITH RECURSIVE keyword in SQL:1999 extended the ability of relational database systems to handle fixed-point computations, unlocking significant performance advantages by allowing computation to move closer to the data. Yet with recursion, SQL becomes a Turing-complete programming language and, with that, unrecoverable safety and correctness risks.
SQL itself lacks a fixed semantics, as the SQL specification is written in natural language, full of ambiguities that database vendors resolve in divergent ways. As a result, reasoning about the correctness of recursive SQL programs must rely on isolated mathematical properties of queries rather than wrestling a unified formal model out of a language with notoriously inconsistent semantics. To address these challenges, we propose a calculus, , that automatically derives mathematical properties from embedded recursive queries and, depending on the database backend, rejects queries that may lead to the three classes of recursive query errors — database errors, incorrect results, and non-termination. We introduce TyQL, a practical implementation in Scala for safe, recursive language-integrated query. Using Named-Tuples and type-level pattern matching, TyQL ensures query portability and safety, showing no performance penalty compared to raw SQL strings while unlocking a three-orders-of-magnitude speedup over non-recursive SQL queries.
1. Introduction
Relational data management systems (RDBMS) are the most popular option for efficient data storage and analytics. In SQL:1999, support was added for recursive common table expressions (CTEs) using the WITH RECURSIVE keyword, enabling fixed-point workloads to execute within RDBMS. Recursive SQL unlocks significant performance gain by pushing computation closer to the data: without WITH RECURSIVE, users must either iterate within the application, use procedural extensions to SQL, or rely on domain-specific backends such as Datalog engines. Repeated round-trips between the application and database or context switches between procedural and plain SQL have significant overhead compared to a single recursive query (Hirn and Grust, 2021). Ideally, applications with fixed-point computations would be able to take advantage of the billions of dollars invested in optimizing relational systems, accessing cutting-edge research and engineering advances without needing to be retrofitted into specialized, domain-specific systems. However, the use of recursion within SQL is widely regarded as “powerful and versatile but notoriously hard to grasp and master” (Hirn and Grust, 2023).
SQL does not have well-established formal semantics as the SQL specification is written in natural language, and is inconsistently followed across RDBMS implementations. Each database vendor adopts its own interpretation, leading to differences in evaluation order, type coercion, null handling, and even fundamental query behavior. As a result, there is no single type-safe or semantically well-defined compilation target for SQL—only a patchwork of best-effort approximations, none widely applied industrially. Yet SQL remains the most popular declarative language due to its powerful optimizability.
Language-integrated query addresses this problem from within the programming language by embedding query constructs using comprehension syntax that supports operations over collections analogous to SQL operations over tables.
Language-integrated query provides safety and portability by allowing users to write their query once and have the language compiler specialize it for multiple backends. The most successful language-integrated query library is the LINQ framework for .NET (Meijer et al., 2006), but the concepts have been applied to many popular languages. Figure 1 shows a classic database query in SQL; canonically in Scala, one of the go-to languages for big data processing and analytics; and our language-integrated query library for Scala, TyQL.
| ⬇ CREATE TABLE Product ( id PRIMARY KEY, name TEXT, price NUMERIC); CREATE TABLE ShipInfo ( id PRIMARY KEY, date DATE); CREATE TABLE Purchase ( id PRIMARY KEY, sid INT, pid INT, cnt INT, total NUMERIC); SELECT pr.name AS name, pu.id AS id, si.date AS sd FROM Purchase pu, Product pr, ShipInfo si WHERE pu.pid = pr.id AND pu.sid = si.id | ⬇ case class Product( id: Int, name: String, price: Double) case class ShipInfo( id: Int, date: LocalDate) case class Purchase( id: Int, sid: Int, pid: Int, cnt: Int, total: Double) def tables = ( prods = Seq[Product](), sInfos = Seq[ShipInfo](), purchs = Seq[Purchase]() ) for purch <- tables.purchs prod <- tables.prods si <- tables.sInfos if purch.pid == prod.id && purch.sid == si.id yield (name = prod.name, id = purch.id, sd = si.date) | ⬇ case class Product( id: Int, name: String, price: Double) case class ShipInfo( id: Int, date: LocalDate) case class Purchase( id: Int, sid: Int, pid: Int, cnt: Int, total: Double) def tables = ( prods = Table[Product](), sInfos = Table[ShipInfo](), purchs = Table[Purchase]() ) for purch <- tables.purchs prod <- tables.prods si <- tables.sInfos if purch.pid == prod.id && purch.sid == si.id yield (name = prod.name, id = purch.id, sd = si.date) |
Classic query comparison.
Ways recursive queries can go wrong.
However, with the addition of WITH RECURSIVE, SQL becomes a Turing-complete programming language and with that comes a host of new correctness and safety concerns orthogonal to classic language-integrated query problems like nested queries or data-type safety. Programming languages have distinct recursion semantics that may only partially align with the SQL specification — and even less with the practical implementations found across modern database systems. The support and implementation of recursive queries vary more widely compared to more well-established SQL features (Winand, 2024). As shown in Figure 2, recursive SQL queries can return incomplete results, throw runtime errors, or never terminate as databases do not effectively identify and prevent the semantic properties of queries that cause problematic behavior. The most reliable way to reason about SQL does not come from a unified semantics imposed across implementations, but rather from the independently applicable and composable mathematical properties of fixed-point computations that underlie relational query evaluation. We propose a solution based on language-integrated query that constrains recursive queries at compile-time using classic mathematical properties, generating only the safe subset of SQL that will not trigger problematic behavior.
To the best of our knowledge, this is the first work that considers the problem of safe language-integrated recursive SQL. We propose holistic compilation of embedded recursive queries to SQL’s WITH RECURSIVE to push computation to the database and enable significant performance benefits. We design the type system to automatically derive relevant mathematical properties of recursive language-integrated queries and reject unsafe queries that risk causing unwanted database behavior. This paper makes the following contributions:
Previous work in language-integrated query did not support recursive SQL. With the addition of a fixed-point operator, SQL becomes a Turing-complete language, leading to new and complex challenges to query correctness and safety. In Section 2.2 we identify the 3 classes of ways that recursive queries “go wrong”: runtime exception, incorrect results, and nontermination, and show what mathematical properties of queries are responsible for each behavior.
We present a calculus, , that automatically checks the mathematical properties responsible for these classes of errors at the host-language compile-time. Sections 4 and 5 show how uses type classes, linear types, and union types to independently encode constraints for each targeted mathematical property so that the generated query is specialized to the semantics of the database backend.
We propose TyQL: practical language-integrated queries in Section 6. TyQL users express their data model using Named-Tuples, enabling efficient implementation within the programming language and straightforward error messages.
In Section 7 we propose a benchmark of recursive queries adapted from industrial benchmarks and academic works in the domains of recursive SQL, Datalog, and Graph Database systems. We conduct a survey of modern relational database systems with respect to query behavior.
We then evaluate TyQL with regard to query coverage and the performance of TyQL with alternative approaches, showing no performance penalty compared to raw SQL strings and a three-orders-of-magnitude speedup over state-of-the-art language-integrated query libraries using non-recursive SQL queries with an in-memory embedded database.
2. Background on Recursive Querying
The algorithms driving modern applications demand increasingly sophisticated query capabilities beyond the limits of traditional select-project-join-aggregate queries. Performance-critical industrial applications, including large-scale program, network, and distributed system analyses, rely on fixed-point computations for data analysis.
2.1. Recursive Query Languages
Datalog is a rule-based declarative query language based in logic programming that has been widely used for recursive query processing, particularly program analysis (Scholz et al., 2016). Datalog programs take the form of a finite set of facts and rules. Datalog rules take the following form:
are predicate symbols, also called relations, and and are terms that can be either variables or constants. The bottom-up evaluation of a Datalog program repeatedly executes unions of conjunctive queries for each derived predicate symbol until an iteration passes where no new facts are discovered (Gottlob et al., 1989). The Semi-Naive evaluation technique only uses facts from the previous iteration to discover new facts (Ullman, 1989).
WITH RECURSIVE was officially added to the SQL standard in 1999 (International Organization for Standardization (1999), ISO). Prior, hierarchical or recursive relationships (for example organizational charts, family trees, graph data, etc.) required use of Datalog, application-side iteration on data extracted from the database, or specialized, non-standard SQL extensions, for example Oracle’s CONNECT BY or PL/SQL. Even after WITH RECURSIVE became standardized, databases have added more powerful or expressive support for recursive queries. For example, DuckDB (duckdb, 2019) allows mutual recursion and MariaDB supports mutual and non-linear recursion (Shalygina and Novikov, 2017). RaSQL (Wang et al., 2020) extends recursive queries to allow certain non-monotone aggregations. Datalog engines, for example Soufflé (Scholz et al., 2016), support expressive forms of recursion like mutual or non-linear recursion but do not implement the SQL standard.
| Recursion | Nonlinear | Mutual | CYCLE | UNION [distinct] | |
| MySQL | ✓(2017) | • (2017) | ✗ | ✗ | ✓ (2017) |
| OracleDB | ✓ (2009) | ✗ | ✗ | ✓ (2009) | ✗ |
| PostgreSQL | ✓ (2009) | ✗ | ✗ | ✓ (2021) | ✓ (2009) |
| SQL Server | ✓ (2005) | • (2008) | ✗ | ✗ | ✗ |
| SQLite | ✓ (2014) | • (2020) | ✗ | ✗ | ✓ (2014) |
| MariaDB | ✓ (2017) | ✓ (2017) | ✓ (2017) | ✓(2020) | ✓ (2017) |
| DuckDB | ✓ (2020) | • (2020) | • (2020) | ✗ (2020) | ✓ (2020) |
Table 1 shows documented recursion support across several modern relational database systems. ✗ indicates that the database does not support the feature and will throw a relevant error message. • indicates syntactic support for a feature, namely, the query will not throw an error but the system may not necessarily have the requisite internal implementation to execute the query successfully. It is possible to run non-linear recursive queries on some databases using syntactic tricks to obfuscate that the query is not linear, e.g. PostgreSQL, but it is not officially supported. ✓ indicates documented support. As illustrated in Table 1, the technical landscape is wide and ever-growing, and no two modern databases support the same set of features. As a result, constraints on recursive queries must be applied independently and specialized to the database.
2.2. How do recursive queries “go wrong”?
In this section, we classify the problems that arise with recursive queries into three areas and identify the mathematical properties responsible for each class of error. The three emergent behaviors targeted are (1) runtime exceptions thrown by the database, (2) incomplete results returned from a query, and (3) nontermination. We show how each targeted behavior is a result of a particular mathematical property, then discuss the cases in which a user may want to relax these restrictions.
2.2.1. Recursive Query Runtime Exception
It is well-known in the Datalog and recursive query literature that to guarantee that the unique and minimal fixed point will be found in a finite number of steps, operations within recursive queries must be monotonic under the ordering of set inclusion: for a monotonic query and databases and , if then (Bancilhon and Ramakrishnan, 1986). Therefore, all non-monotonic operators, i.e. aggregation or negation, must happen between strata of recursive queries to guarantee termination (that is, aggregations can only be applied to the results of recursive queries after they have completed, not to results within the bodies of recursive queries).
Bill-of-materials
Transitive closure
Transitive closure
Examples of non-monotonic, non-linear, and non-set semantic queries.
For example, Figure 3 shows the widely-used Bill-of-Materials problem (Wang et al., 2020) that models items sold by a business that are made out of sub-parts. We use the relation SubParts(part, sub) for each item a business sells and its sub-parts (and sub-part sub-parts, etc.); and the relation BasicParts(part, days) for base parts and how many days it takes for that part to arrive from a supplier. The query waitFor determines when a part will be ready, given that a part will be ready on the day the last subpart arrives. Naively, this query can be represented as Figure 3. However, the presence of the MAX aggregation within the body of the recursive query definition means that the query is not monotonic under the ordering of set inclusion. The SQL standard specifies that only monotonic queries are supported. This constraint is usually checked by the query engine, so the emergent behavior is a run-time error for the application as it must wait for the round-trip time for the query to be sent to the database and the error returned. All widely-used commercial databases and Datalog engines officially support only monotonic operations within recursion, so for the query in Figure 3 to pass a database query compiler check, the MAX aggregation must be moved out of the recursive query. The set of database query errors that are unrelated to recursion, for example datatype errors, are well covered by previous work in language-integrated query that is complementary to our approach (Cheney et al., 2013). Therefore, the mathematical property that leads to recursive query runtime exceptions is non-monotonicity.
However, recent advances in recursive query engines have shown that some forms of aggregation (Wang et al., 2020; Abo Khamis et al., 2022) are permissible within recursion without losing termination guarantees. While this has not yet been implemented widely in commercial systems, it shows when a user may still want to “turn off” the monotonicity constraint for certain backends.
2.2.2. Incomplete Results
The query to compute the transitive closure of a directed graph can be expressed in Datalog with two rules: and . This Datalog program is equivalent to the SQL query in Figure 3. However, given a 3-step input graph, i.e. {(0, 1), (1, 2), (2, 3)} this query returns a result with 5 rows (the input edges plus {(0, 2), (1, 3)}) in PostgreSQL (v15) and DuckDB (v1.1). However, when the equivalent query is run in Datalog, the result has 6 rows (the previous results plus (0, 3)). Quickly stepping through the graph shows that 3 is reachable from 0, so why is it not in the RDBMS result?
Unlike Datalog, SQL requires that all WITH RECURSIVE queries be linearly recursive. The way this restriction is often enforced is to limit references to the recursive relation to only once within the recursive subquery. However this is a purely syntactic restriction, so simple aliasing can easily prevent this check from happening, while other databases do not check query linearity at all. Around the time that SQL’99 was preparing to add recursive queries to SQL, there was a belief that “most ‘real life’ recursive queries are indeed linear” (Bancilhon and Ramakrishnan, 1986). While this belief is no longer widely held, this assumption is built into the implementation of many database systems. Within Datalog engines that use the Semi-Naive evaluation algorithm (Ullman, 1989), for the transitive closure example, at each iteration the “new” data is computed with the subquery expressed in the unnamed relational algebra: where represents the data discovered in the previous iteration, and represents all the data discovered so far over the course of iteration. However, if a query is guaranteed to be linear, then the query can be optimized to read only , so the previous query becomes which may cause the algorithm to terminate too early for non-linear queries. Therefore, the mathematical property that leads to incomplete results is non-linearity.
As there are modern databases that do not perform any checks that queries are linear and instead allow the query to return incorrect results, it is of utmost importance that users be prevented from unknowingly sending queries that will silently return invalid results. However recently, MariaDB claimed to add support for non-linear queries (Shalygina and Novikov, 2017), so a user may wish to turn off this constraint depending on their database backend.
2.2.3. Nontermination
WITH RECURSIVE expands the expressive power of SQL to be Turing-complete, so it is possible to represent infinite loops and nontermination. Assuming the SQL specification is followed, that is, all queries are monotonic and linear, what causes nontermination?
Given the query is monotonic over set inclusion, nontermination is the result of infinitely growing intermediate relations. Datalog queries are guaranteed to generate only finite relations, given they are correctly range-restricted over a finite domain (Gottlob et al., 1989). There are a few key differences in the semantics of RDBMS compared to Datalog engines that cause the mathematical properties that provide termination guarantees to be violated.
Range-restriction is defined as the constraint that all variables in the head of a Datalog rule must appear in the body of the rule. Range-restriction applied to SQL queries requires the project clause of each recursive query to contain only constants or references to columns present in relations in the FROM clause. However SQL queries are not limited to a finite domain, as column-level “constructor” operations like addition or string concatenation can arbitrarily introduce new elements. The Datalog literature refers to queries that do not contain operations that can infinitely grow the domain as having the constructor-freedom property (Arntzenius and Krishnaswami, 2016). Finally, the bottom-up evaluation of logic programs assumes set semantics, which provides termination guarantees at the cost of a performance penalty (checking for and removing duplicates). However, RDBMS use bag semantics unless otherwise specified, so a reachability query using bag semantics over data containing cycles will repeatedly re-discover the same paths. Only certain operators like DISTINCT or UNION will use set semantics. For example, given a linearly recursive transitive closure query, and a cyclic input graph, i.e. {(0, 1), (1, 0)}, the query in Figure 3 will infinitely recur. Therefore, the mathematical properties that provide termination guarantees are constructor-freedom and set-semantics.
However, not all queries require set semantics to terminate, although given only the query, it is undecidable whether a query will terminate when using less restrictive duplicate checks (Maher and Ramakrishnan, 1989). If the data does not contain cycles, then the user may prefer to use bag semantics to avoid wasted time on duplicate elimination. Other RDBMS require the use of UNION ALL between the base and recursive case of a recursive query, for example SQLServer and OracleDB, and provide alternative language structures to check for infinite recursion (e.g. CYCLE in OracleDB and PostgreSQL v14+). Users may also wish to turn off the constructor-freedom constraint, as not all constructors lead to infinitely growing domains. Range-restriction, however, will always be enforced. Given the high penalty of an infinitely recurring query, both on the application and the database itself, as other users may also see reduced performance from cross-query interference, it is important to give users the ability to prevent infinitely recurring queries, even at the cost of possible performance overhead or limits to expressibility.
3. Overview of Embedded Safe Recursion
RestrictedExpr and RestrictedQuery types and the connection to each query property.
In Section 2.2, we established the connection between certain problematic database behavior and the mathematical properties of (1) range-restriction, (2) monotonicity, (3) linearity, (4) set-semantics, and (5) constructor-freedom. To avoid generating SQL queries that cause this unwanted behavior, we must design our system to reject queries that do not hold these properties independently.
Figure 4 illustrates how the types representing relations and queries (RestrictedQuery[A, C, D]), and expressions on columns (RestrictedExpr[A, S]) track these properties. Each query is built by applying combinators like map and flatMap on relations. The RestrictedQuery type represents relations that are recursively defined, the A type parameter represents the column-type of the relation, the C type represents the category, e.g. if a query uses bag semantics and can contain duplicates, the D type represents the set of dependency edges between recursive relations. The RestrictedExpr type represents expressions that operate over columns of recursive relations, with A representing the result type and S representing the shape of the expression, e.g. if it contains any non-monotonic operations. Section 4 shows how our system captures these semantic properties of queries and Section 5 uses the host-language type system to translate them into syntactic checks.
The goal of TyQL is not to provide evaluation semantics for recursive queries because database backends vary widely in their interpretation of recursive queries, but there is one consistency across all evaluation engines: the precedence graph of relation dependencies. Therefore, we design our type system to target properties on this graph. Intuitively, each of the five properties can be understood as properties on the program graph of the recursive query. Relations are represented with nodes, dependencies on source relations are represented as directed edges, and recursive queries are subgraphs with cycles. We provide an abstract interpretation that maps terms in to a graph representation in the appendix.
4. Enforcing Mathematical Properties
with
In this section, we present the calculus for recursive queries. We show how the mathematical properties of monotonicity, linearity, set-semantics, range-restriction, and constructor-freedom are encoded into the type system independently so that the restrictions can be composed depending on the targeted database system. Because RDBMS have fragmented support for recursion, yet offer significant performance benefits, it is essential that the type system allow restrictions to be tuned and not overly constrict real-world queries. We first present the DSL type system in Section 4 and then show how to extend it to a 2-level DSL and host language type system in Section 5.
4.1. Syntax and Base Type System
Syntax of Types and Terms
BOM query
The non-recursive part of extends T-LINQ (Cheney et al., 2013). The syntax of types and terms and the typing judgments are presented in Figure 5. ranges over variables. The column types represent database column types, Named-Tuples represent row types, represents a query that returns a collection of rows of type A, and represents an aggregation that returns a single scalar result of type A. Tuples and function types are used to construct the combinators () that represent database operations. The combinators follow the precedent set by T-LINQ, and we add to represent database-level recursion. We refer to “column-level expressions” as any expression that would go into the filter or project subclause of a SQL query: for map, it would be SELECT a + 1 FROM table vs. for aggregate it would be SELECT sum(a) + 1 FROM TABLE. Both a + 1 and sum(a) + 1 are column-level expressions. Query-level expressions refer to the entire query expression and operations that take relations, e.g. union.
Following the convention set by T-LINQ, we assume a signature that maps each constant c and operator to the typing rule. is useful to abstract over the large set of operations that share the same typing behavior. The operations stored in () can be column-level expressions (, e.g. ); column-level aggregations (, e.g. sum), or relation-level expressions (, e.g. union). Examples of operations stored in are presented in Figure 5.
DSL Typing rules with only range-restricted recursion
We build our syntax on the Nested Relational Algebra extended with a “bounded fixpoint” operator (Suciu, 1997). To illustrate an example, in Figure 2 we show a recursive query on the BOM database from Section 2.2 in SQL and . In the TyQL implementation, for-comprehensions desugar to combinators but we do not include for-comprehensions in to keep the syntax minimal. In SQL, a recursive query is composed of a “base case” that contains no references to recursive relations, and a “recursive case” where references to the recursive relation are allowed. Although some databases do not syntactically require a base case, we enforce this in because recursion without a base case leads to undefined behavior. Figure 7 shows the typing rules for the DSL, which closely follows the precedent type system in T-LINQ, with the addition of fix. Judgment states that term has type in the DSL type environment . The fix function defines recursive relations within a single stratum. For , fix takes as arguments a tuple of “base case” queries and a function . The function takes as arguments a tuple of references to the recursive relations being defined and returns a tuple of “recursive case” definitions . Each is composed from the recursive references (and any relations in scope that are defined outside the body of fix) using the combinators in . In the next section, we show how we apply the type system to independently identify and prevent the mathematical properties identified in Section 2.2.
4.2. Range-restricted Recursion
Range-restriction is the constraint that every variable appearing in the head of a rule appears in the body of that rule, to avoid rules that yield infinite relations. Because this constraint does not depend on the semantics of the database system, it is encoded into the basic FIX-D typing rule in Figure 7 by enforcing the recursive references and the recursive definitions all have the same type, and that is a Named-Tuple. As Named-Tuples must have the same key and value types, order, and arity to be considered the same type, this restriction enforces that all variables in the head of a the rule (e.g. columns in project) are present in the body of the rule (e.g. recursive definitions returned by ).
Definition 1.
A function such that holds the range-restricted property if .
4.3. Monotone Recursion
| Syntax | |||||
Operations that are non-monotonic under set inclusion are defined as scalar operations. For example, the query SELECT max(a) + 1 FROM TABLE is non-monotonic and produces a single value, while SELECT a + 1 FROM table is monotonic and nonscalar. Both scalar and nonscalar expressions can be constructed from each other: + is nonscalar while sum is scalar but the expressions sum(a + 1) or sum(a) + 1 are both valid and should produce an expression of type scalar, while SELECT x + (SELECT max(y) FROM table1) FROM table2 should produce an expression of type nonscalar (as the subquery does not change that the full expression is monotonic). The type system must be able to distinguish the shape of the entire expression (no matter how nested) so that well-formed terms of type Query are guaranteed to return nonscalar results and are monotonic under set inclusion. The required changes to the type system to restrict recursive queries to monotone operations are shown in Figure 8. We add a type to wrap expressions on columns so that can track the shape () of the expression. We split the OP-D rule into separate rules for , , and . By EXPR-OP-D and AGG-OP-D, any expression that contains an arbitrarily nested sub-expression with type parameter will be of type (via Shape). For example:
The expression on the left contains a nested scalar operation (sum) so the resulting type will be , while the expression on the right contains only nonscalar expressions, so even with the same operator () the result type is The MAP-D and FLATMAP-D rules are refined to only accept functions that contain non-scalar expressions, while AGGREGATE-D and GROUPBY-D only accept functions that contain at least one scalar sub-expression. With these restrictions, expressions of type Query in are guaranteed to contain only monotonic operations.
The monotonicity property is relevant only between the relations in a single stratum (Bancilhon and Ramakrishnan, 1986). To limit the monotonicity restriction to only recursive references, we introduce a RestrictedQuery type and refine FIX-D so that the arguments to must be of type RestrictedQuery. Crucially, RestrictedQuery can only be derived by calling combinators on the arguments passed to the function given to fix as it is not available in the syntax. We update the MAP-D, FLATMAP-D, FILTER-D, and REL-OP-D so if they are passed any arguments of type RestrictedQuery the result type will be RestrictedQuery.
To reduce the number of rules in Figure 8, we define a meta-syntax helper to abstract away the differences between rules for combinators that take RestrictedQuery and Query. Without , we could equivalently define 4 rules for flatMap: if and , the result type will be , however if either or are RestrictedQuery the result will be . As AGGREGATE-D and GROUPBY-D are valid only on Query types, will reject non-monotonic operations on recursive references.
Definition 2.
A function such that holds the monotonicity property if .
4.4. Linear Recursion
fix restricted to be linear
We approach the problem of encoding linear recursion in the type system with a linear type with respect to recursive references. To accomplish this, we split the problem into two sub-constraints: an affine check (every relation is never used more than once) and a relevant check (all relations are used in at least one recursive definition). The affine check is per-relation, while the relevant check is for all relations, as not every relation needs to use every other relation within a single stratum. The extensions to the type system to restrict queries to linear recursion are shown in Figure 9.
A type parameter of type is added to to model the dependencies of each query. Uniqueness of references is enforced by using the argument position within the fix function: for a recursive function , will be a tuple containing the constant integer type , and will contain a tuple with the constant integer type .
Some database backends, for example DuckDB, allow recursive queries nested within outer queries to return columns from the outer query (which can itself be recursive, e.g. fix within a fix). However, this would break linearity, so must be able to differentiate between references per call to fix. Each reference is tagged with a singleton type that is unique for each fix invocation. References are considered equal only if both the and the constant integer type are the same. This ensures that if fix is called within the scope of references from another fix function, the inner fix cannot return terms derived from the outer function’s parameters. For example:
For the affine check, the dependencies of each recursive definition must contain no duplicates. For the relevant check, all parameters must appear at least once in all . This is implemented by using the set of constant integer types from 1 to the number of arguments to fix decorated with to check that all elements are present at least once in all the s () and never more than once in each ().
The REL-OP-D rule is updated so if any arguments are RecursiveQuery, the result will be a RecursiveQuery that collates (only) the dependencies of the recursive relations. Similarly to Figure 8, to reduce the number of rules in Figure 9 we define a meta-syntax helper to abstract away the differences between rules for combinators that take RestrictedQuery, Query, or both.
Definition 3.
If then holds the linearity property if it holds both the relevant and affine properties. The relevant property holds if and the affine property if .
4.5. Set semantic Recursion
| Syntax | |||||
fix restricted to be set semantic
To restrict fix to allow only set semantics within recursive queries, we track the semantics of each operation in a type parameter, similarly to how we tracked monotonicity. We update so that all relation-level operations reflect their semantics. The required changes to the syntax and typing rules are presented in Figure 10. For example:
Definition 4.
A function such that holds the set-semantic property if .
4.6. Constructor-Freedom
Constructors, e.g. column-level operations that produce new values and therefore grow the program domain, are represented with the types stored in and . To ensure that an expression is constructor-free, it must not contain any of those operations. We apply this restriction in by modifying the functions accepted by map, flatMap, and filter. The typing rules are updated in a similar way to the monotonicity check: arguments and return types of in MAP-D, FLATMAP-D, and FILTER-D must be of type RestrictedExpr so by construction, and that produce new values cannot be applied to RestrictedExpr. This is by far the most restrictive property, and similar systems like Soufflé (souffle, 2019) have chosen to allow limited unsoundness in exchange for expressiveness, a well-established approach to avoiding excessively compromising usability (Livshits et al., 2015). Because each restriction in is independently applicable, users are free to choose if they wish to enforce it.
Definition 5.
A function such that holds the constructor-freedom property if and is of type RestrictedExpr.
4.7. Nesting in
The problem of deeply nested datatypes and query avalanches is well covered by previous work in NRC and T-LINQ. The unnesting and normalization approach taken by T-LINQ is applicable to non-recursive queries, or solely within the bodies of recursive queries, but it is not applicable between non-recursive and recursive subqueries because the fixpoint introduces a strict evaluation boundary—often referred to as a ‘pipeline breaker’, which prevents the same normalization approach. Therefore, we have chosen to simplify the syntax of to restrict nested column types, although in our implementation (discussed in Section 6) nested datatypes are supported.
For nested recursion (i.e., fix in fix), as described earlier, some database systems such as DuckDB allow nested recursive queries to return columns from the outer query. This can be only problematic when it violates linearity, which is handled by the type system as explained in Section 4.4.
5. with Host Language Embedding
| ScalarelseNonScalar |
Host Typing rules with all restrictions
The power of language-integrated query comes from the tight embedding of the DSL with the general-purpose host language. The DSL environment represents staged computation: all functions in the DSL syntax serve to construct a type-level AST but do not actually execute any database operations. To generate queries that are run on a database and returned to an application written in a general-purpose language, we need to extend to be an embedded DSL in a host language. In Figure 11, based on the precedent set by T-LINQ, we extend with a second type environment for host terms, , and update our DSL typing rules accordingly. Judgment states that host term has type in type environment , and judgment states that quoted term has type in host type environment and DSL type environment .
We reuse the syntax of types and terms in Figure 5 for the host language, with a few key additions. DSL types are represented in the host language by wrapping them in Expr types, so if a DSL expression represents a database row of type then the type of this expression in the host language will be . Translation functions from host terms to DSL terms are added: toExpr converts base host-language types to Expr of base types and toRow converts Named-Tuples of Expr to Exprs of Named-Tuple. Finally, an execution function run is added to execute the query expressed in the DSL and returns a List of the query result type to the application. TO-EXPR, TO-ROW, RUN-QUERY, and RUN-AGG rules define the types for toExpr, toRow, and run. The typing rules for in Figure 11 combine all the restrictions presented in Figures 8, 9, 10 with the addition of rules to handle interactions between the DSL and host language. The typing rule for the fully-restricted host language fix is shown in Figure LABEL:fig:fix-impl.
The DSL typing judgments must also be updated slightly. Each judgment in the DSL type system operates under both and and an additional rule, LIFT, defines the connection between the host typing environment and the DSL typing environment.
6. Implementation
In this section, we describe the implementation of TyQL, our type-safe embedded query library based on . Figure LABEL:fig:fix-impl shows the implementation of fix in TyQL with all restrictions applied and the typing rule from with all restrictions applied, renamed to match the variable names in TyQL for readability. We include the side-by-side comparison to emphasize that our implementation and our theory are closely knit. The architecture of TyQL is summarized in Figure 12.
6.1. Modeling Rows in Scala
Representing database rows in statically typed languages like Scala is challenging because the compiler must allow on-the-fly composition of types without losing the advantages of static typing. Beyond type safety, static typing also enables powerful IDE features like code completion. This is nontrivial because operations like join and project take collections of rows and produce new collections that may be of a completely different structure, but still need to support element access and be type-checked when used later in the program.
Type computations in JVM languages cannot create classes, so it is impossible to dynamically generate Scala’s classes. Structural types allow abstraction over existing classes but generally require reflection or other mechanisms to support dynamic element access. Named-Tuples, released in Scala 3.6.0, are represented within the Scala compiler as pairs of Tuples, where names are stored as a tuple of constant strings and the values are stored in regular Scala tuples. Tuples are preferrable over case classes because tuples are lightweight structures that avoid additional JVM object allocation and dispatch overhead. In contrast, classes generate separate class definitions at compile-time, increasing memory usage and execution overhead and lead to bloated bytecode generation. By using tuples, TyQL benefits from more compact and efficient runtime representations, reducing both memory footprint and execution latency.
Named-Tuples are ordered, which gives them an advantage over structural types for modeling rows because there is (1) better integration with the language since they share the same representation as regular tuples, and (2) efficient and natural traversal order allows the formulation of type-generic algorithms. Because Named-Tuples can be decomposed into head *: tail, they have an induction principle and can be iterated over generically and without the use of an auxiliary data structure like a dictionary. Method overloading can catch common mistakes like using map instead of flatMap and provide useful, domain-specific error messages. For example, a simple type error generates the following error messages in TyQL and in a state-of-the-art query library ScalaSQL:
Architecture of TyQL
6.2. Type-Level ASTs and Constraints
TyQL maintains a hierarchy of query representations, illustrated in Figure 13. Rows are represented as Named-Tuples, AST expressions as structural types that wrap row types, and entire queries or tables as DatabaseASTs. These can be of type Query, which allows chaining of further relation-level operations, or Aggregation, which represents a scalar result from the database. Aggregations are also subtypes of expressions, as many databases allow aggregations to be nested within queries at the expression level. Because aggregations and non-scalar expressions must share a supertype, monotonicity of expressions is tracked in a type member ExprShape that can be either Scalar or NonScalar. The category of the result, either bag or set, is tracked with Category. RestrictedQuery wraps Query but does not extend it.
6.2.1. Selection
The Expr class represents AST expressions and extends Scala’s Selectable trait. This means that element accesses are syntactic sugar for a call to the method selectDynamic, which maps field names to values. So calling x.t where x: Expr[A] and A: (t: T) will return a new AST node Select[T](x, "t"). Implicit conversions (or explicit if the user prefers) are used to lift native Scala datatypes to AST expressions.
6.2.2. Projection
The method that models projection is defined in Figure 14. Type-level pattern matching with Match Types (Blanvillain et al., 2022) is used to lift tuples-of-expressions to expressions-of-tuples (lines 1-3) and to enforce that all values of the tuple are of type Expr (lines 4-5). The IsTupleOfExpr constraint first converts the Named-Tuple into a regular Tuple with NamedTuple.DropNames, then takes the Union type of all values in the tuple with Tuple.Union, and finally constrains the resulting Union type to be a subtype of a non-scalar Expr with <:<. The toRow method is defined (lines 6-7) on Named-Tuples of type Expr and constructs an instance of the Project AST node. This method is used as an implicit conversion between Named-Tuple-of-Expr and Project[A] (line 8). TyQL reifies the constant string keys in Named-Tuples using type tags (ResultTag) so that the generated queries are more readable, for example t.map(r => (maxCount = max(r.count)) generates the SQL query SELECT MAX(count) as maxCount FROM t.
Hierarchy of types in TyQL.
Projection implementation.
6.2.3. Join
Join operations are represented by FlatMap AST nodes, subtypes of DatabaseAST as they are relation-level operations. Each FlatMap, Map, or Aggregate has a source subtree of type DatabaseAST and a Fun subtree that represents function application. Unrolling of nested Fun nodes is done during query generation due to simpler implementation, but could also happen at the type-level with intricate use of Match Types.
6.2.4. GroupBy
We design GroupBy in TyQL slightly differently than to allow for incrementally chaining map, groupBy, and having. Due to SQL semantics, groupBy and having should operate over the type of the source relations, not the result type of the preceding expression. This is a challenge, especially when the preceding expression is a join, as there will be multiple input relations. The way this is addressed in TyQL is by tracking the types of the source relations of compound statements at the type-level.
6.2.5. Recursive Constraints
Query syntax trees are built and verified by the host language compiler, and constraints are enforced with type classes and implicit evidence. As shown in Figure LABEL:fig:fix-impl, each of the implicit constraints in the definition of fix in TyQL have a corresponding premise in (with the exception of line #7, which is essentially a hint to the compiler to understand the structure of the dependency tuple, DT). TyQL uses the annotation @implicitNotFound to customize error messages, so that the user can see which constraint failed: i.e., if their recursive query was affine but not relevant. For systems that do not support mutual recursion, we do not have to include this constraint in the type system because we can simply limit the number of arguments to the fix function to a single relation, which prevents the construction of mutually recursive queries. Unique query references are implemented using constant integer types and unique types are implemented using anonymous classes and value types.
Figure LABEL:fig:fix-impl shows the fully restricted fix, but the restrictions can be made configurable: for example, if a user wants a fix function with all restrictions except set-semantics, they simply call val bagFixFn = fix((ConstructorFree(), Monotone(), BagResult(), Linear())). This is implemented by adding additional type parameters for each restriction [RCF <: ConstructorFreedom, RM <: MonotoneRestriction, RC <: ResultCategory, RL <: LinearRestriction] to line 2 and an additional argument to line 3: (options: (RCF, RM, RC, RL)). The type arguments are forwarded to the match types ToRestrictedQueryRef and ToRestrictedQuery in lines 4 and 8 that will evaluate to different types depending on those arguments.
6.3. Query Generation
To allow the composition of sub-queries and sub-expressions as well as abstraction over DSL expressions, TyQL generates SQL lazily. Users can peek at the generated SQL using the helper method toSQLString, but must call run to execute an expression. When it is time to generate SQL, TyQL constructs a tree-based intermediate representation that closely resembles a database query plan. Comprehension syntax naturally lends itself to nested queries, and significant work has been done on the problem of unnesting comprehension-generated queries into flat SQL queries (Cheney et al., 2014). TyQL merges composable sequences of operations. For example, table.filter(_.x > 1).filter(_.y < 2) can be merged into table.filter(r => r.x > 1 && r.y < 2). A similar approach is applied to sequences of flatMap, for example the query t1.flatMap(r1 => t2.flatMap(r2 => t3.map(r3 => (r1 = r1.x, r2 = r2.x, r3 = r3.x)))) is flattened into a single JOIN node (as opposed to a binary, left-deep tree).
7. Evaluation
| ⬇ // TyQL: val base = Table[(dst:Int, cst:Int)]("base") val edges = Table[(src:Int, dst:Int, cst:Int)]("edge") base.fix(sp => edges.flatMap(edge => sp.filter(s => s.dst == edge.src) .map(s => (dst = edge.dst, cst = s.cst + edge.cst)) .distinct) .aggregate(s => (dst = s.dst, cst = min(s.cst))) .groupBy(s => (dst = s._1.dst)) | ⬇ -- SQL: CREATE TABLE edge (src INT, dst INT, cst INT); CREATE TABLE base (dst INT, cst INT); WITH RECURSIVE path AS ( (SELECT * FROM base) UNION (SELECT edge.dst, path.cst + edge.cst FROM edge, path WHERE path.dst = edge.src)) SELECT path.dst, MIN(path.cst) FROM path GROUP BY path.dst |
SSSP Query.
| Query | Description | Linear | Mono-tone | Set | Mut-ual | CF |
| Even-Odd | ◼ Mutual recurive program generates even/odd numbers. | ✓ | ✓ | ✓ | ✓ | ✗ |
| CSPA | ⚫ Graspan’s Context Sensitive Pointer Analysis, static analysis for identifying aliases. (Wang et al., 2017) | ✗ | ✓ | ✓ | ✓ | ✓ |
| Company Control (CC) | ▲ Calculates the complex controlling relationships between companies. (Mumick et al., 1990) | ✓ | ✗ | ✓ | ✓ | ✗ |
| PointsToCount (PTC) | ⚫ Find the count of objects that a variable of a given name may point to (Abeysinghe et al., 2024) (Java variant). | ✗ | • | ✓ | ✓ | ✓ |
| Chain Of Trust (COT) | ◼ Security query where two entities trust each other if they are direct friends or if their friends trust each other. | ✓ | ✓ | ✗ | ✓ | ✓ |
| Java Points To (JPT) | ⚫ Field-sensitive subset-based points to analysis for an object-oriented language. (Madsen et al., 2016) | ✗ | ✓ | ✗ | ✓ | ✓ |
| Party | ★ Mutually recursive social media graph algorithm.(Wang et al., 2020) | ✓ | ✗ | ✗ | ✓ | ✗ |
| CBA | ◼ Constraint-based analysis query. (souffle, 2019) | ✗ | • | ✗ | ✓ | ✓ |
| Single Source Shortest Path (SSSP) | ★ Computes the shortest path from a given source node to all other nodes in a weighted graph. (Wu et al., 2022) | ✓ | • | ✓ | ✗ | ✗ |
| Ancestry | ▲ Find descendants of a person who are of the same generation. (Bancilhon and Ramakrishnan, 1986) | ✓ | ✓ | ✓ | ✗ | ✗ |
| Andersen’s Points To (APT) | ⚫ Context-insensitive, flow-insensitive interprocedural pointer analysis. (Fan et al., 2019a) | ✗ | ✓ | ✓ | ✗ | ✓ |
| All Pairs Shortest Path (APSP) | ★ Compute the shortest paths between all pairs of nodes in a weighted graph. (Wu et al., 2022) | ✗ | ✗ | ✓ | ✗ | ✗ |
| Graphalytics (TC) | ★ Directed cyclic graph reachability query. Uses list data structure to check for cycles. (Angles et al., 2020) | ✓ | ✓ | ✗ | ✗ | ✗ |
| Bill of Materials (BOM) | ▲ Business query for how many days it would take to deliver a product made of subparts. (ibm, 2021) | ✓ | • | ✗ | ✗ | ✓ |
| Orbits | ◼ Orbits of cosmological objects. (souffle, 2019) | ✗ | • | ✗ | ✗ | ✓ |
| Data Flow | ◼ Models control flow through read/write/jump instructions. (souffle, 2019) | ✗ | ✓ | ✗ | ✗ | ✓ |
Linear, Set, Mutual, CF (constructor-free): ✓ query has the property.
Monotone: ✓ no aggregation • stratified aggregation ✗ unstratified aggregation.
In this section, we evaluate TyQL with respect to real-world use cases and how effectively TyQL can constrain recursive queries without compromising on usability. To illustrate the range of real-world queries, we introduce a recursive query benchmark (RQB) of 16 queries taken from open source code repositories and recent publications from various query domains, including business management, program analysis, graph queries, and other classic fixed-point problems. We give a survey of modern relational database systems using our benchmark, identifying the ways each query can go wrong and which RDBMS supports each query property or combination of properties.
We use the RQB to evaluate TyQL with respect to query coverage: how many queries that would go wrong are prevented from compiling, and impact: if TyQL were to allow the query to compile, how it would fail on the database. Lastly, we evaluate TyQL with respect to performance and compare to alternative approaches.
7.1. Recursive Query Benchmark and Survey of RDBMS
In this section, we present a recursive query benchmark (RQB) comprising 16 representative queries across diverse domains such as business management, program analysis, graph queries, and classic fixed-point problems and explore how each query behaves on different RDBMS.
The goal of our benchmark is to simulate a broad range of recursive queries. We have identified five discrete properties of recursive queries that have the greatest impact on behavior: (1) linearity, (2) monotonicity, (3) set semantics, (4) mutual recursion, and (5) constructor-freedom. All queries are range-restricted. We have selected 16 queries to represent classes of queries for each combination of properties 1-4. Since constructor-freedom applies independently to any query, we included a mix of queries with this property. Identifying the frequency of each query across real-world applications remains future work. All queries run and terminate on at least one of the evaluated database systems. For the monotonicity property, we have selected a mix of queries without aggregation or negation, with stratified-aggregation, and with unstratified aggregation. Table 2 illustrates the benchmark property matrix. As an example, we illustrate the Single Source Shortest Path query expressed in TyQL and in SQL in Figure 15. The full set of queries in the recursive benchmark expressed in both TyQL and the generated SQL are included in the appendix.
| DuckDB | Postgres | SQLite | MariaDB | What restriction is violated? | |
| Even-Odd | ✓ | ✗ | ✗ | ◗ | mutual, cf |
| CSPA | ◗ | ✗ | ✗ | ✗ | linear, mutual |
| CC | ◗ | ✗ | ✗ | ✗ | monotone, mutual, cf |
| PTC | ◗ | ✗ | ✗ | ✗ | linear, mutual |
| COT | ✓ ❘ ▼✳ | ✗ | ✗ | ✗ | set, mutual |
| JPT | ◗ ❘ ▼✳ | ✗ | ✗ | ✗ | linear, set, mutual |
| Party | ◗ | ✗ | ✗ | ▼ | monotone, set, mutual, cf |
| CBA | ◗ ❘ ▼✳ | ✗ | ✗ | ✗ | linear, set, mutual |
| SSSP | ✓ ❘ ▼✳ | ✓ ❘ ▼ | ✓ ❘ ▼ | ✓ | cf |
| Ancestry | ✓ | ✓ | ✓ | ✓ | cf |
| APT | ◗ | ◗ | ✗ | ◗ | linear |
| APSP | ◗ | ◗ | ✗ | ✓ | linear, monotone, cf |
| TC | ✓ | ✓ | ✓ | ✓ | set, cf |
| BOM | ✓ ❘ ▼✳ | ✓ ❘ ▼✳ | ✓ ❘ ▼✳ | ✓ | set |
| Orbits | ◗ ❘ ▼✳ | ◗ ❘ ▼✳ | ✗ | ▼ | linear, set |
| Data Flow | ◗ ❘ ▼✳ | ◗ ❘ ▼✳ | ✗ | ▼ | linear, set |
As shown in Section 2.2, support for recursive queries varies widely across RDBMS. To get a sense of what classes of queries each system supports, we ran the 16 queries with cyclic input data and with acyclic input data. The results of the RQB on four RDBMS are presented in Table 3. We represent queries that terminated with the full result with ✓, database error with ✗, incomplete results with ◗, and nontermination with ▼. If the query exhibited different behavior based on if the input data was cyclic or not, for example if the query terminates correctly with acylic data but not cyclic data, then the query is represented as ✓ ❘ ▼ with the acyclic data on the left and cyclic data on the right. We consider “incomplete” results to be data that is missing results that would have been present using a non-SQL version of each algorithm, either via general-purpose imperative programs or Datalog programs. The • and ▼ classification indicates that it is possible to find an input dataset that returns incomplete results or does not terminate. The databases used are DuckDB v1.1, Postgres v15, SQLite v3.39, and MariaDB v11.5.2 with the configuration ---skip-standard-compliant-cte.
7.2. TyQL Coverage and Impact
In this section, we use the RQB to evaluate the effectiveness of TyQL in achieving comprehensive query coverage while preventing queries that may go wrong from compiling, illustrated in Table 3.
The Ancestry query exemplifies the class of queries that exhibit only the properties fully supported by the SQL specification: it is linear, monotonic, set-semantic and does not contain any mutual recursion. Only with constructor-freedom enforced will TyQL reject this query. We provide this option to show the completeness of our approach, although we expect most real-world queries to fall into this category and do not anticipate that it will be widely applied because it is very restrictive for real-world use cases.
For the queries that return incorrect results, all are either mutually recursive or non-linear. The missing tuples are those that would have been generated by intermediate results from previous iterations that are “forgotten” by the SQL engine due to the implementation only reading tuples derived in the immediately preceding iteration. For the queries that do not terminate, all but one use bag semantics. Careless use of a bag-semantic query can cause nontermination for input data sets that have cycles (✳ next to the non-terminating queries indicates that if we change our bag semantics to set semantics, then the query will terminate on both acyclic and cyclic input data). Users may prefer to pay a performance penalty due to the duplicate elimination cost of set semantics to avoid the risk that their queries will not terminate. Nonterminating queries can have a significant performance impact, both on the application and on the other users of the database, due to interference. Some databases allow you to set a max recursion depth to avoid infinite recursion, although it is not obvious how deep to set this without trial and error. It is clear that the impact of incorrect results is more damaging than a database throwing an error. However, if nontermination is more impactful than incomplete results depends on the context of the application and database system. So far, the only relational data management system we have seen to officially include non-linear or mutual recursion is MariaDB (Shalygina and Novikov, 2017), although in our experiments we found that it often returned slightly different results from Datalog engines.
The TC query returns correctly on all evaluated systems even though it uses bag semantics, even with cyclic input data. The reason for this is that the query itself checks for duplicates: in DuckDB, the query has a list that tracks visited nodes, while in systems that do not support lists, the query appends to a string for each node it has visited. Conversely, the SSSP query will not terminate on cyclic data even with set semantics. The reason for this is the cost propagation, where each tuple generated at each recursive step includes the “weight” of the newly discovered path. If the query reaches a cycle in the data, the weight of the path will be infinitely increasing and the ever-changing cost column will prevent the set difference from removing already-discovered paths. The property responsible for this is constructor-freedom: the TC query constructs new values used to detect cycles even using bag semantics while the SSSP query constructs new values that lead to cyclic reinforcement and non-termination.
In summary, we evaluate the ability of TyQL to identify queries that will fail and find that it can be tuned to successfully reject all problematic queries. As with many type systems, TyQL takes a conservative approach to correctness and can reject queries that may, if the data has certain properties, return successfully. To maximize usability and practicality, TyQL users always have the choice to tune how restrictive the fix operator is, potentially allowing the full RQB to compile.
7.3. Performance and State-of-the-Art
| Query | Size | It | TyQL (s) | Collections (s) | ScalaSQL (s) | SQL String (s) | vs. Collections | vs. ScalaSQL | vs. SQL String |
| Ancestry | 0.01MB | 3 | 0.008 | 0.002 | 0.047 | 0.005 | 0.25X* | 6.12X | 0.70X* |
| 10MB | 15 | 0.209 | 3.712 | 0.423 | 0.214 | 17.73X | 2.02X | 1.02X* | |
| 100MB | 189 | 39.469 | TO | TO | 39.522 | ¿15.20X | ¿15.20X | 1.00X* | |
| APT | 0.01MB | 3 | 0.013 | 0.001 | 0.055 | 0.012 | 0.11X | 4.17X | 0.91X* |
| 0.02MB | 4 | 0.020 | 4.859 | 0.087 | 0.016 | 241.38X | 4.30X | 0.82X* | |
| 0.04MB | 9 | 0.038 | 58.665 | 0.225 | 0.035 | 1553.75X | 5.96X | 0.94X* | |
| ASPS | 0.01MB | 3 | 0.012 | 0.003 | 0.044 | 0.011 | 0.25X* | 3.52X | 0.88X* |
| 1MB | 3 | 0.048 | 33.197 | 0.188 | 0.045 | 694.97X | 3.93X | 0.95X* | |
| 5MB | 4 | 0.177 | TO | 0.670 | 0.172 | ¿3389.83X | 3.78X | 0.97X* | |
| BOM | 0.01MB | 2 | 0.009 | 0.002 | 0.036 | 0.009 | 0.24X* | 4.02X | 0.96X* |
| 2MB | 5 | 0.083 | 139.448 | 0.565 | 0.075 | 1675.25X | 6.79X | 0.90X* | |
| 20MB | 22 | 1.108 | TO | 14.777 | 1.115 | ¿541.52X | 13.34X | 1.01X* | |
| CBA | 0.02MB | 9 | 0.028 | 0.006 | 0.311 | 0.025 | 0.22X | 11.03X | 0.88X* |
| 0.1MB | 1 | 0.020 | TO | 0.054 | 0.017 | ¿30000.00X | 2.73X | 0.83X* | |
| 0.2MB | 1 | 0.024 | TO | 0.059 | 0.020 | ¿25000.00X | 2.49X | 0.83X* | |
| CC | 0.01MB | 3 | 0.011 | 0.002 | 0.070 | 0.009 | 0.21X | 6.19X | 0.82X* |
| 1MB | 3 | 0.295 | 13.839 | 0.766 | 0.300 | 46.91X | 2.60X | 1.02X* | |
| 1.5MB | 3 | 0.831 | 40.566 | 1.016 | 0.788 | 48.79X | 1.22X | 0.95X* | |
| CSPA | 0.01MB | 5 | 0.024 | 0.003 | 0.171 | 0.018 | 0.12X | 7.18X | 0.76X* |
| 2MB | 11 | 1.069 | TO | 5.712 | 1.109 | ¿561.27X | 5.34X | 1.04X* | |
| 10MB | 14 | 31.675 | TO | 318.531 | 31.571 | ¿18.94X | 10.06X | 1.00X* | |
| Data Flow | 0.01MB | 3 | 0.006 | 0.002 | 0.048 | 0.005 | 0.31X | 7.39X | 0.81X* |
| 0.03MB | 3 | 0.010 | 17.675 | 0.058 | 0.008 | 1768.88X | 5.84X | 0.84X* | |
| 0.05MB | 5 | 0.014 | 257.405 | 0.129 | 0.013 | 18981.25X | 9.52X | 0.93X* | |
| Even-Odd | 0.01MB | 17 | 0.014 | 0.004 | 0.327 | 0.009 | 0.29X | 23.56X | 0.62X* |
| 1MB | - | 181.488 | TO | TO | 178.274 | ¿3.31X | ¿3.31X | 0.98X* | |
| 2MB | - | 450.077 | TO | TO | 454.288 | ¿1.33X | ¿1.33X | 1.01X* | |
| JPT | 0.02MB | 3 | 0.014 | 0.003 | 0.077 | 0.012 | 0.18X | 5.37X | 0.85X* |
| 0.05MB | 16 | 0.055 | TO | 0.835 | 0.053 | ¿10909.09X | 15.27X | 0.96X* | |
| 0.1MB | - | 0.510 | TO | TO | 0.493 | ¿1176.47X | ¿1176.47X | 0.97X* | |
| Orbits | 0.01MB | 2 | 0.012 | 0.002 | 0.038 | 0.010 | 0.17X* | 3.17X | 0.80X* |
| 1MB | 2 | 0.077 | TO | 0.392 | 0.075 | ¿7792.21X | 5.11X | 0.98X* | |
| 10MB | 2 | 0.432 | TO | 0.744 | 0.383 | ¿1388.89X | 1.72X | 0.89X* | |
| Party | 0.01MB | 5 | 0.012 | 0.004 | 0.099 | 0.010 | 0.32X* | 8.40X | 0.83X* |
| 2MB | 5 | 0.093 | TO | 0.954 | 0.089 | ¿6451.61X | 10.23X | 0.95X* | |
| 20MB | 7 | 0.858 | TO | 4.796 | 0.873 | ¿699.30X | 5.59X | 1.02X* | |
| PTC | 0.02MB | 3 | 0.016 | 0.003 | 0.085 | 0.012 | 0.16X | 5.26X | 0.77X* |
| 0.05MB | 16 | 0.052 | TO | 0.841 | 0.048 | ¿11538.46X | 16.13X | 0.91X* | |
| 0.1MB | - | 2.516 | TO | TO | 2.451 | ¿238.47X | ¿238.47X | 0.97X* | |
| SSSP | 0.01MB | 5 | 0.009 | 0.003 | 0.068 | 0.009 | 0.29X* | 8.01X | 1.01X* |
| 10MB | 9 | 0.026 | 0.578 | 0.152 | 0.026 | 21.84X | 5.75X | 0.97X* | |
| 25MB | 42 | 0.163 | 233.752 | 1.773 | 0.167 | 1435.36X | 10.89X | 1.02X* | |
| TC | 0.01MB | 2 | 0.007 | 0.001 | 0.031 | 0.005 | 0.21X | 4.48X | 0.73X* |
| 5MB | 5 | 0.021 | 0.333 | 0.079 | 0.017 | 15.90X | 3.79X | 0.83X* | |
| 10MB | 10 | 0.064 | 6.634 | 0.250 | 0.062 | 103.29X | 3.90X | 0.97X* | |
| COT | 0.01MB | 7 | 0.011 | 0.003 | 0.160 | 0.008 | 0.25X | 14.00X | 0.72X* |
| 1MB | 7 | 0.124 | TO | 2.611 | 0.124 | ¿4838.71X | 21.03X | 1.00X* | |
| 15MB | - | 1.324 | TO | 9.976 | 1.356 | ¿453.17X | 7.54X | 1.02X* |
In this section we evaluate TyQL with respect to performance. Application developers who wish to run fixed-point algorithms on data stored in a RDBMS have several options. The most immediate choice is to simply read data into memory and then execute their algorithm using the programming language constructs. Unfortunately, the Scala Collections API does not include a built-in fixed-point operator, so users must implement their own iterative control flow. The benefit of this approach is that it is fully customizable, yet it puts the burden onto the developer and the machine where the application is running. Alternatively, users may offload computation to the database. A natural way to do this would be to use language-integrated query to compose queries that are compiled to SQL and sent to the database. Yet if the query library does not support WITH RECURSIVE, then users will need to handle control flow at the application level and send only non-recursive queries at each iteration, or default to raw SQL strings.
7.3.1. Experimental Setup
Table 4 shows the execution time (s) of the RQB presented in Section 7.1. The “Collections” column shows the execution time of the query implemented purely within the programming language using the Collections API and no database backend. The “ScalaSQL” column shows the execution time of the latest state-of-the-art language-integrated query library ScalaSQL, using non-recursive SQL queries (as recursion is not supported). This approach is representative of other language-integrated query libraries in Scala that support only non-recursive SQL. Because we use an embedded database system that runs within the same process as the application, avoiding the overhead of round-trips between database and application, this approach is equivalent to a PL/SQL approach. The “SQL String” column shows the execution time of sending raw strings directly to the JDBC driver without any language-integration. The “TyQL” column shows the execution time of the query using TyQL. The final columns, “vs. Collections”, “vs. ScalaSQL”, and “vs. SQL String” show the speedup of TyQL over each respective approach. In the speedup columns, “¿” indicates that the baseline did not terminate within a 10-minute timeout so we calculate the minimum speedup. The “It” column states the number of iterations needed for the Collections API and non-recursive SQL to reach a fixed point (it was not possible to extract the number of iterations from the DuckDB internals without impacting the result) and the “size” column states the total size of the input relations. * indicates the benchmarking harness margin of error exceeds the difference in execution time (max margin of error - difference average execution time) ¿ 0 and can be considered to have equivalent runtimes due to normal variation in the JVM JIT.
The fixpoint implementation used in both the non-recursive SQL and the collections-only implementation is tail-recursive and based on the canonical example given in Scala by Example (Odersky, 2005) extended to use the same bottom-up Semi-Naive evaluation algorithm used internally in the database. The database used is in-memory DuckDB v1.1 with JDBC driver v.1.1.0 and queries that risked nontermination are run with set-semantics. Each query is run on synthetic data of three different input relation sizes. Experiments are run on an Intel(R) Xeon(R) Gold 5118 CPU @ 2.30GHz (2 x 12-core) with 395GB RAM, on Ubuntu 22.04 LTS with Linux kernel 6.5.0-17-generic and Scala 3.5.1-RC1 (using experimental) with JDK 17.0.9, OpenJDK 64-Bit Server VM, 17.0.9+9-jvmci-23.0-b22, with -Xmx8G. We use Java Benchmarking Harness (JMH) (Shipilev et al., 2022) v1.37 with -i 5 -wi 5.
7.3.2. Experimental Results and Analysis
| Slowdown vs. SQL String | Risks | ||||||
| Max | Min | Avg | Data-type errors | Runtime DB error | Incorrect results | Nontermination | |
| SQL String | 1X | 1X | 1X | ✗ | ✗ | ✗ | ✗ |
| Collections | 30000X | 0.12X | 2787.12X | ✓ | - | ✗ | ✗ |
| ScalaSQL | 1176.47X | 1.29X | 37.27X | ✓ | - | ✗ | ✗ |
| TyQL | 1X | 1X | 1X | ✓ | ✓ | ✓ | ✓ |
The smallest of the three input data sets shows the range of data sizes for which the Collections API outperforms the other approaches due to avoiding the overhead of database connection and initialization. ScalaSQL has a higher overhead than TyQL or SQL strings, as the query initialization overhead happens at every iteration. The medium-sized dataset shows data input sizes where the ScalaSQL approach outperforms the Collections API due to several reasons: the data management system query optimizer can effectively select efficient query plans and join algorithms; as the data is not sent back to the application at each iteration, overheads due to boxing and unboxing of primitive types are avoided until iteration has concluded. For this input data size TyQL outperforms the other approaches due to avoiding multi-query overhead, storing and copying intermediate relations, and internal database optimizations. The memory usage of Collections is higher than either the TyQL or ScalaSQL because the data is stored externally, and does not put any pressure on the JVM garbage collector. The largest data size shows the cases where the Collections API and non-recursive SQL approaches may run out of memory or time-out after 10-minutes. The graph algorithm queries (Ancestry, TC, SSSP) are run on larger datasets, while program analysis queries (APT, CBA, Data Flow, JPT, PTC) are run on smaller datasets to avoid all systems running out of memory.
Statistics over the data presented in Table 4 are shown in Table 5. Queries whose JMH error margin exceeds any difference in execution time are considered equivalent (indicated as 1X). TyQL shows no performance penalty compared to raw SQL strings, with a significant performance gain over Collections and ScalaSQL using non-recurive SQL.
The options present a classic tradeoff between customizability and performance: raw SQL strings show the best performance but provide no safety guarantees, hand-written imperative programs provide expressibility and flexibility, language-integrated non-recursive queries put less burden on the developer and show better performance, and lastly TyQL puts the least burden on the developer and shows performance equivalent to raw SQL strings while providing safety guarantees.
8. Related Work
8.1. Embedded Query Languages
Type-safe embedded query languages using collections was pioneered by (Buneman et al., 1995) and Kleisli (Wong, 2000) and found commercial success in LINQ, formalized in T-LINQ (Cheney et al., 2013). Recently, there has been renewed interest in extending language-integrated query beyond core SQL, for example temporal queries (Fowler et al., 2022) and privacy-aware distributed queries (Salvaneschi et al., 2019). While neither LINQ nor these systems target recursive SQL, there has been work in general-purpose functional languages with fixed-point semantics (Madsen et al., 2016; Pacak and Erdweg, 2022) that can operate as a functional Datalog. TyQL shares the goal of using functional abstractions to structure recursion while ensuring safety through a well-defined type system. These approaches extend Datalog semantics while TyQL specifically targets RDBMS, which requires abstracting over different SQL semantics to ensure backend portability.
There has been significant interest in embedded SQL support in Scala. ScalaQL (Spiewak and Zhao, 2010) uses anonymous inner classes to model row types, while Slick (slick, 2011; Shaikhha, 2013) provides SQL embedding in Scala with two frontends using macros (Jovanovic et al., 2014) and the implicit resolution in Scala’s type system. ScalaSQL (scalasql, 2024) uses higher-kinded case classes to model rows, and Quill (quill, 2016) uses refinement types, macros and quotation to compile SQL queries at Scala compile-time. Most of these libraries aim to provide ergonomic SQL APIs that expose the SQL query and data model to the user, that is, they take the spirit of the Collections API while still exposing the SQL query model to users. In this work we aim for transparent persistence (Wiedermann and Cook, 2007) so the distinction between processing of data stored in the native language collections or a database is as minimal as possible.
8.2. Recursion and Relational Databases
Researchers have attempted to address the impedence mismatch problem within the data management system: object-oriented or document databases provide data models and query languages that integrate cleanly with general-purpose programming languages but are more difficult to optimize for efficient execution (Ramakrishnan and Gehrke, 2003). Alternatively, object-relational mapping libraries (ORMs) attempt to provide object-oriented abstractions on top of relational databases but can also suffer from performance penalties and obscure query behavior (Goldschmidt et al., 2008).
The problem of extracting relational algebraic properties from general-purpose programs is known as the query extraction problem and has been successfully applied to synthesize queries from application code (Ramachandra et al., 2018). Recent work in this area has used SQL WITH RECURSIVE as a compilation target when compiling user-defined functions written in procedural language extensions like PL/SQL (Duta et al., 2019) or Python functions (Fischer et al., 2022). The aim of this line of work is to accept arbitrary programs written in general-purpose languages and compile them to SQL, while the goal of TyQL is to provide type-safe recursive language-integrated query using a compile-time-restricted embedded DSL. RDD2SQL (Zhang et al., 2023) uses counterexample-guided inductive synthesis to automatically translate functional database APIs like Spark RDDs into SQL but does not specifically target recursion. Novel extensions to SQL with cleaner recursion semantics have been proposed (Hirn and Grust, 2023) but are not widely implemented in commercial databases.
Datalog is one of the most successful query languages with recursion capabilities, which has been used for fixpoint computations in program analysis (Scholz et al., 2016; Fan et al., 2019b; Jordan et al., 2016). As the core Datalog disallows non-monotone operations, various extensions of it have been proposed (Klopp et al., 2024b; Shaikhha et al., 2024; Klopp et al., 2024a; Wang et al., 2021). Flix (Madsen et al., 2016; Madsen and Lhoták, 2020) is a general-purpose language that supports Datalog queries as first-class citizens. Datafun (Arntzenius and Krishnaswami, 2016) is a functional Datalog variant that tracks monotonicity, constructor-freedom, and range-restriction via type-level constraints. Unlike TyQL, Flix and Datafun have their own runtime and execution engine based on Datalog semantics, while TyQL translates into a subset of safe SQL.
9. Conclusion
Recursive queries are difficult to use and support across databases is fragmented and chaotic, yet the performance gains from offloading computation to the database can be massive. To make recursion easier and safer, without preventing users from expressing real-world queries, we propose TyQL, a language-integrated recursive query in Scala, a popular, practical programming language. TyQL provides a clean abstraction for recursive queries while ensuring correctness and safety at compile-time that is specialized to the database. We formalized the constraints with , which prevents unwanted behavior across databases, including errors, incorrect results, and nontermination.

### v2

@@
 Abstract.
-Performance-critical industrial applications, including large-scale program, network, and distributed system analyses, rely on fixed-point computations.
-The introduction of recursive common table expressions (CTEs) using the WITH RECURSIVE keyword in SQL:1999 extended the ability of relational database systems to handle fixed-point computations, unlocking significant performance advantages by allowing computation to move closer to the data.
-Yet with recursion, SQL becomes a Turing-complete programming language and, with that, unrecoverable safety and correctness risks.
-SQL itself lacks a fixed semantics, as the SQL specification is written in natural language, full of ambiguities that database vendors resolve in divergent ways.
-As a result, reasoning about the correctness of recursive SQL programs must rely on isolated mathematical properties of queries rather than wrestling a unified formal model out of a language with notoriously inconsistent semantics.
-To address these challenges, we propose a calculus, , that automatically derives mathematical properties from embedded recursive queries and, depending on the database backend, rejects queries that may lead to the three classes of recursive query errors — database errors, incorrect results, and non-termination.
+Performance-critical applications, including large-scale program analyses, graph analyses, and distributed system analyses, rely on fixed-point computations.
+The introduction of recursive queries using the WITH RECURSIVE keyword in SQL:1999 extended the ability of relational database systems to handle fixed-point computations, unlocking significant performance advantages by allowing computation to move closer to the data.
+Yet, with the addition of recursion, SQL becomes a Turing-complete programming language with new correctness and safety risks.
+SQL lacks a fixed semantics, as the SQL specification is written in natural language, full of ambiguities that database vendors resolve in divergent ways.
+As a result, reasoning about the correctness of recursive SQL programs must rely on isolated properties of queries rather than wrestling a unified formal model out of a language with notoriously inconsistent semantics.
+To address these challenges, we propose a calculus, , that derives properties from embedded recursive queries using the host-language type system and, depending on the database backend, rejects queries that may lead to the three classes of recursive query errors - runtime database exceptions, incorrect results, and non-termination.
 We introduce TyQL, a practical implementation in Scala for safe, recursive language-integrated query.
-Using Named-Tuples and type-level pattern matching, TyQL ensures query portability and safety, showing no performance penalty compared to raw SQL strings while unlocking a three-orders-of-magnitude speedup over non-recursive SQL queries.
+TyQL uses modern type system features of Scala 3, namely Named-Tuples and type-level pattern matching, to ensure query portability and safety.
+Furthermore, TyQL shows no performance penalty compared to raw SQL strings while unlocking a three-orders-of-magnitude speedup over non-recursive SQL queries.
 1. Introduction
-Relational data management systems (RDBMS) are the most popular option for efficient data storage and analytics.
-In SQL:1999, support was added for recursive common table expressions (CTEs) using the WITH RECURSIVE keyword, enabling fixed-point workloads to execute within RDBMS.
-Recursive SQL unlocks significant performance gain by pushing computation closer to the data: without WITH RECURSIVE, users must either iterate within the application, use procedural extensions to SQL, or rely on domain-specific backends such as Datalog engines.
-Repeated round-trips between the application and database or context switches between procedural and plain SQL have significant overhead compared to a single recursive query (Hirn and Grust, 2021).
-Ideally, applications with fixed-point computations would be able to take advantage of the billions of dollars invested in optimizing relational systems, accessing cutting-edge research and engineering advances without needing to be retrofitted into specialized, domain-specific systems.
-However, the use of recursion within SQL is widely regarded as “powerful and versatile but notoriously hard to grasp and master” (Hirn and Grust, 2023).
-SQL does not have well-established formal semantics as the SQL specification is written in natural language, and is inconsistently followed across RDBMS implementations.
-Each database vendor adopts its own interpretation, leading to differences in evaluation order, type coercion, null handling, and even fundamental query behavior.
-As a result, there is no single type-safe or semantically well-defined compilation target for SQL—only a patchwork of best-effort approximations, none widely applied industrially.
-Yet SQL remains the most popular declarative language due to its powerful optimizability.
-Language-integrated query addresses this problem from within the programming language by embedding query constructs using comprehension syntax that supports operations over collections analogous to SQL operations over tables.
-Language-integrated query provides safety and portability by allowing users to write their query once and have the language compiler specialize it for multiple backends.
-The most successful language-integrated query library is the LINQ framework for .NET (Meijer et al., 2006), but the concepts have been applied to many popular languages.
-Figure 1 shows a classic database query in SQL; canonically in Scala, one of the go-to languages for big data processing and analytics; and our language-integrated query library for Scala, TyQL.
-| ⬇ CREATE TABLE Product ( id PRIMARY KEY, name TEXT, price NUMERIC); CREATE TABLE ShipInfo ( id PRIMARY KEY, date DATE); CREATE TABLE Purchase ( id PRIMARY KEY, sid INT, pid INT, cnt INT, total NUMERIC); SELECT pr.name AS name, pu.id AS id, si.date AS sd FROM Purchase pu, Product pr, ShipInfo si WHERE pu.pid = pr.id AND pu.sid = si.id | ⬇ case class Product( id: Int, name: String, price: Double) case class ShipInfo( id: Int, date: LocalDate) case class Purchase( id: Int, sid: Int, pid: Int, cnt: Int, total: Double) def tables = ( prods = Seq[Product](), sInfos = Seq[ShipInfo](), purchs = Seq[Purchase]() ) for purch <- tables.purchs prod <- tables.prods si <- tables.sInfos if purch.pid == prod.id && purch.sid == si.id yield (name = prod.name, id = purch.id, sd = si.date) | ⬇ case class Product( id: Int, name: String, price: Double) case class ShipInfo( id: Int, date: LocalDate) case class Purchase( id: Int, sid: Int, pid: Int, cnt: Int, total: Double) def tables = ( prods = Table[Product](), sInfos = Table[ShipInfo](), purchs = Table[Purchase]() ) for purch <- tables.purchs prod <- tables.prods si <- tables.sInfos if purch.pid == prod.id && purch.sid == si.id yield (name = prod.name, id = purch.id, sd = si.date) |
-Classic query comparison.
+Fixed-point workloads contain computations that apply a query repeatedly, over the result of the previous iteration, until the result no longer changes.
+SQL:1999 (International Organization for Standardization, 1999) introduced the WITH RECURSIVE keyword, enabling fixed-point workloads to be executed within relational database management systems (RDBMS).
+Recursive SQL unlocks significant performance gain, in part by pushing computation closer to the data (Hirn and Grust, 2021).
+Ideally, applications with fixed-point computations would be able to leverage the billions of dollars invested in optimizing RDBMS, accessing cutting-edge research and engineering advances without needing to be retrofitted into specialized, domain-specific systems, such as Datalog engines.
+Despite excellent performance, the use of recursion within SQL is widely regarded as “powerful and versatile but notoriously hard to grasp and master” (Hirn and Grust, 2023).
+The difficulty is exacerbated by the lack of a well-established formal semantics, as the SQL specification is written in natural language and is inconsistently followed across RDBMS implementations.
+Each database vendor adopts its own interpretation, leading to differences in evaluation order, type coercion, null handling, and even fundamental query behavior (Neumann and Leis, 2024).
+As a result, there is no single type-safe or semantically well-defined compilation target for SQL, only a patchwork of best-effort approximations that either make simplifying assumptions or target a limited subset of SQL (Negri et al., 1991; Guagliardo and Libkin, 2017; Veanes et al., 2010; Chu et al., 2017; Malecha et al., 2010; Ricciotti and Cheney, 2022), none widely applied industrially.
+Language-integrated query addresses inconsistencies across RDBMS from within the programming language, providing safety and portability by allowing users to write their query once and have the language compiler type-check and specialize it for multiple backends.
+The most successful language-integrated query library is the LINQ framework for .NET (Meijer et al., 2006), and the concepts have been widely adopted across modern programming languages and databases (slick, 2011; quill, 2016; jslinq, 2025; diesel, 2025).
+Yet neither LINQ nor other language-integrated libraries support recursion.
+With the addition of WITH RECURSIVE, SQL becomes a Turing-complete programming language, and with that comes a host of new concerns orthogonal to classic language-integrated query problems like nested queries or data-type safety.
 Ways recursive queries can go wrong.
-However, with the addition of WITH RECURSIVE, SQL becomes a Turing-complete programming language and with that comes a host of new correctness and safety concerns orthogonal to classic language-integrated query problems like nested queries or data-type safety.
-Programming languages have distinct recursion semantics that may only partially align with the SQL specification — and even less with the practical implementations found across modern database systems.
-The support and implementation of recursive queries vary more widely compared to more well-established SQL features (Winand, 2024).
-As shown in Figure 2, recursive SQL queries can return incomplete results, throw runtime errors, or never terminate as databases do not effectively identify and prevent the semantic properties of queries that cause problematic behavior.
-The most reliable way to reason about SQL does not come from a unified semantics imposed across implementations, but rather from the independently applicable and composable mathematical properties of fixed-point computations that underlie relational query evaluation.
-We propose a solution based on language-integrated query that constrains recursive queries at compile-time using classic mathematical properties, generating only the safe subset of SQL that will not trigger problematic behavior.
-To the best of our knowledge, this is the first work that considers the problem of safe language-integrated recursive SQL.
-We propose holistic compilation of embedded recursive queries to SQL’s WITH RECURSIVE to push computation to the database and enable significant performance benefits.
-We design the type system to automatically derive relevant mathematical properties of recursive language-integrated queries and reject unsafe queries that risk causing unwanted database behavior.
+Figure 1 shows how recursive SQL queries can throw runtime errors, return incorrect results, or never terminate.
+Furthermore, the support and implementation of recursive queries vary more widely across commercial databases compared to other SQL features (Winand, 2024).
+As a result, the most effective and practical way to reason about recursive SQL does not come from a unified semantics imposed across database implementations.
+The key insight behind our paper is to provide a DBMS-agnostic reasoning framework over recursive queries based on independently applicable and composable mathematical properties of fixed-point computations that determine how queries will behave: range-restriction, monotonicity, mutual-recursion, linearity, set-semantics, and constructor-freedom.
+We propose a solution based on language-integrated query that constrains queries at application compile-time, enabling significant performance benefits while generating only the safe subset of SQL that will not trigger runtime exceptions, nontermination, or incorrect results.
+Figure 2 shows a classic recursive database query for the transitive closure of a directed graph in our language-integrated query library for Scala, TyQL, and in SQL.
+We design the type system to automatically derive the six relevant mathematical properties and, depending on the database backend, reject unsafe queries.
 This paper makes the following contributions:
 Previous work in language-integrated query did not support recursive SQL.
-With the addition of a fixed-point operator, SQL becomes a Turing-complete language, leading to new and complex challenges to query correctness and safety.
-In Section 2.2 we identify the 3 classes of ways that recursive queries “go wrong”: runtime exception, incorrect results, and nontermination, and show what mathematical properties of queries are responsible for each behavior.
-We present a calculus, , that automatically checks the mathematical properties responsible for these classes of errors at the host-language compile-time.
-Sections 4 and 5 show how uses type classes, linear types, and union types to independently encode constraints for each targeted mathematical property so that the generated query is specialized to the semantics of the database backend.
+With the addition of a fixed-point operator, SQL becomes a Turing-complete programming language, leading to new and complex challenges to query correctness and safety.
+In Section 2 we provide background on recursive database queries and in Section 3 we identify the three classes of ways that recursive queries “go wrong”: runtime exception, incorrect results, and nontermination, and show the mathematical properties of queries that are responsible for each behavior.
+We present a calculus, , that automatically checks the six properties responsible for these classes of errors at the host-language compile-time.
+Sections 5 and 6 show how uses type classes, linear types, and union types to independently encode constraints for each targeted query property so that the generated query is specialized to the semantics of the database backend.
 We propose TyQL: practical language-integrated queries in Section 7.
@@
 We then evaluate TyQL with regard to query coverage and the performance of TyQL with alternative approaches, showing no performance penalty compared to raw SQL strings and a three-orders-of-magnitude speedup over state-of-the-art language-integrated query libraries using non-recursive SQL queries with an in-memory embedded database.
-2. Background on Recursive Querying
-The algorithms driving modern applications demand increasingly sophisticated query capabilities beyond the limits of traditional select-project-join-aggregate queries.
-Performance-critical industrial applications, including large-scale program, network, and distributed system analyses, rely on fixed-point computations for data analysis.
-2.1. Recursive Query Languages
-Datalog is a rule-based declarative query language based in logic programming that has been widely used for recursive query processing, particularly program analysis (Scholz et al., 2016).
-Datalog programs take the form of a finite set of facts and rules.
-Datalog rules take the following form:
+Examples of Transitive Closure Query in Recursive SQL, and TyQL.
+2. Background on Recursion in Databases
+Modern applications demand increasingly sophisticated query capabilities beyond the limits of traditional select-project-join-aggregate queries.
+Performance-critical industrial applications, including large-scale program analyses, network analyses, artificial intelligence, and distributed system analyses, rely on fixed-point computations (Sahebolamri et al., 2022; Backes et al., 2019; Das et al., 2019; Darwiche, 2020; Bellomarini et al., 2018; Vistrup, 2025).
+Datalog (Gottlob et al., 1989) is a recursive query language based on logic programming that has been particularly useful for program analysis (Scholz et al., 2016).
+A Datalog program is a finite set of facts and rules where rules take the form: .
 are predicate symbols, also called relations, and and are terms that can be either variables or constants.
@@
 The Semi-Naive evaluation technique only uses facts from the previous iteration to discover new facts (Ullman, 1989).
-WITH RECURSIVE was officially added to the SQL standard in 1999 (International Organization for Standardization (1999), ISO).
-Prior, hierarchical or recursive relationships (for example organizational charts, family trees, graph data, etc.) required use of Datalog, application-side iteration on data extracted from the database, or specialized, non-standard SQL extensions, for example Oracle’s CONNECT BY or PL/SQL.
-Even after WITH RECURSIVE became standardized, databases have added more powerful or expressive support for recursive queries.
-For example, DuckDB (duckdb, 2019) allows mutual recursion and MariaDB supports mutual and non-linear recursion (Shalygina and Novikov, 2017).
-RaSQL (Wang et al., 2020) extends recursive queries to allow certain non-monotone aggregations.
-Datalog engines, for example Soufflé (Scholz et al., 2016), support expressive forms of recursion like mutual or non-linear recursion but do not implement the SQL standard.
-| Recursion | Nonlinear | Mutual | CYCLE | UNION [distinct] | |
-| MySQL | ✓(2017) | • (2017) | ✗ | ✗ | ✓ (2017) |
-| OracleDB | ✓ (2009) | ✗ | ✗ | ✓ (2009) | ✗ |
-| PostgreSQL | ✓ (2009) | ✗ | ✗ | ✓ (2021) | ✓ (2009) |
-| SQL Server | ✓ (2005) | • (2008) | ✗ | ✗ | ✗ |
-| SQLite | ✓ (2014) | • (2020) | ✗ | ✗ | ✓ (2014) |
-| MariaDB | ✓ (2017) | ✓ (2017) | ✓ (2017) | ✓(2020) | ✓ (2017) |
-| DuckDB | ✓ (2020) | • (2020) | • (2020) | ✗ (2020) | ✓ (2020) |
+WITH RECURSIVE was added to the SQL standard in 1999 (International Organization for Standardization, 1999), based on the same theoretical foundation and Semi-Naive evaluation technique developed for Datalog.
+Prior, hierarchical or recursive relationships (e.g., organizational charts, family trees, graph data, etc.) required use of Datalog, application-side iteration on data extracted from the database, or specialized, non-standard SQL extensions, for example Oracle’s CONNECT BY keyword or procedural extensions like PL/SQL.
+It has been shown that repeated round-trips between the application and database or context switches between procedural and plain SQL have significant overhead compared to a single recursive query (Hirn and Grust, 2021).
+Recursive queries define recursive relations, and are composed of a base case query and a recursive case query that references the recursive relations.
+For example, the query in Figure 2 defines a recursive relation named ”path”, the base case is the ”edges” relation, and the recursive case joins the ”edges” and ”path” relations, producing the transitive closure of all nodes.
+Since 1999, databases have added even more powerful or expressive support for recursion.
 Table 1 shows documented recursion support across several modern relational database systems.
@@
 • indicates syntactic support for a feature, namely, the query will not throw an error but the system may not necessarily have the requisite internal implementation to execute the query successfully.
-It is possible to run non-linear recursive queries on some databases using syntactic tricks to obfuscate that the query is not linear, e.g.
-PostgreSQL, but it is not officially supported.
-✓ indicates documented support.
+| WITH | Range-Rest | Agg | Mutual | Non-linear | Set | Const-Free | |
+| RECURSIVE | (P1) | (P2) | (P3) | (P4) | (P5) | (P6) | |
+| MySQL | ✓ (2017) | ✓ (2017) | ✗ | ✗ | • (2017) | ✓ (2017) | ✓ (2017) |
+| OracleDB | ✓ (2009) | ✓ (2009) | ✗ | ✗ | ✗ | ✗ | ✓ (2009) |
+| PostgreSQL | ✓ (2009) | ✓ (2009) | ✗ | ✗ | ✗ | ✓ (2009) | ✓ (2009) |
+| SQL Server | ✓ (2005) | ✓ (2005) | ✗ | ✗ | • (2008) | ✗ | ✓ (2005) |
+| SQLite | ✓ (2014) | ✓ (2014) | ✗ | ✗ | • (2020) | ✓ (2014) | ✓ (2014) |
+| MariaDB | ✓ (2017) | ✓ (2017) | ✗ | ✓ (2017) | ✓ (2017) | ✓ (2017) | ✓ (2017) |
+| DuckDB | ✓ (2020) | ✓ (2020) | ✓ (2020) | • (2020) | • (2020) | ✓ (2020) | ✓ (2020) |
+Range-restriction.
+The “Range-Rest” column in Table 1 indicates queries must be range-restricted.
+The Datalog literature defines the property of range-restriction (Gottlob et al., 1989) as the constraint that variables on the left-hand-side of a Datalog rule must appear in the right-hand-side.
+Applied to recursive SQL queries, this constraint requires the project clause to contain only constants or references to columns present in relations in the FROM clause.
+All RDBMS require range-restriction.
+Monotonicity.
+The “Agg” column indicates support for aggregation operations within the bodies of recursive queries.
+So far, the only commercial database to allow for aggregations within recursive queries is DuckDB, but academic systems like RaSQL (Wang et al., 2020) extend recursive queries to allow certain restricted aggregations.
+Aggregations are restricted because it is well-known in the Datalog and recursive query literature that to guarantee that the unique and minimal fixed point will be found in a finite number of steps, operations within recursive queries must be monotonic under the ordering of set inclusion.
+A monotonic query is defined as a query and databases and , such that if then (Bancilhon and Ramakrishnan, 1986).
+A query is monotonic if adding more data to its input does not remove data from its output, and aggregations like MAX or NOT EXISTS can violate this property.
+Datalog variants with aggregation generally require stratification, where recursive relations are grouped into layers, or strata, where each layer can only depend on the results of earlier layers.
+To ensure termination, aggregation is allowed across strata, but not within a single stratum.
+Recursive SQL enforces stratification by construction, as each WITH RECURSIVE query is defined as one stratum.
+Because aggregation is allowed between recursive queries, all the commercial RDBMS presented in Table 1 except DuckDB support stratified aggregation, while DuckDB is more expressive than stratified Datalog.
+Mutual-recursion.
+The “Mutual” column indicates support for mutually recursive queries, where two or more recursive relations refer to each other in a cyclic dependency.
+Support for mutual recursion is essential for expressing many static analyses or bidirectional graph traversals, and is built into Datalog semantics but not SQL.
+Recently, MariaDB added support for mutual recursion, and DuckDB inlines relations such that it is possible to express some mutually recursive queries.
+Linearity.
+The “Non-linear” column indicates support for non-linear recursive queries, in which a recursive relation is referenced more than once within a recursive subquery.
+Non-linear queries are particularly useful for program analysis queries (Wang et al., 2017; Madsen et al., 2016; Abeysinghe et al., 2024).
+DuckDB does not check queries for linearity, PostgreSQL applies a simple and easily avoided syntactic check, and MariaDB supports non-linear queries.
+Set-semantics.
+The “Set” column indicates support for the UNION operator to combine the base and recursive cases, which uses set semantics.
+Without UNION, the default is UNION ALL, which uses bag (i.e., multiset) semantics.
+Some RDBMS, e.g., OracleDB, require the use of bag semantics.
+Constructor-freedom.
+The “Const-Free” column indicates support for queries that violate the constructor-freedom property.
+The Datalog literature refers to queries that do not contain interpreted functions over infinite domains (e.g., integer arithmetic) as having the constructor-freedom property (Arntzenius and Krishnaswami, 2016).
+All RDBMS support constructors.
+Pure Datalog does not support constructors or aggregations.
+Modern Datalog engines, for example Soufflé (Scholz et al., 2016), support expressive forms of recursion like mutual or non-linear recursion and use set semantics but miss out on the wide adoption and significant engineering resources that go into making commercial RDBMS efficient.
 As illustrated in Table 1, the technical landscape is wide and ever-growing, and no two modern databases support the same set of features.
-As a result, constraints on recursive queries must be applied independently and specialized to the database.
-2.2. How do recursive queries “go wrong”?
-In this section, we classify the problems that arise with recursive queries into three areas and identify the mathematical properties responsible for each class of error.
-The three emergent behaviors targeted are (1) runtime exceptions thrown by the database, (2) incomplete results returned from a query, and (3) nontermination.
-We show how each targeted behavior is a result of a particular mathematical property, then discuss the cases in which a user may want to relax these restrictions.
-2.2.1. Recursive Query Runtime Exception
-It is well-known in the Datalog and recursive query literature that to guarantee that the unique and minimal fixed point will be found in a finite number of steps, operations within recursive queries must be monotonic under the ordering of set inclusion: for a monotonic query and databases and , if then (Bancilhon and Ramakrishnan, 1986).
-Therefore, all non-monotonic operators, i.e.
-aggregation or negation, must happen between strata of recursive queries to guarantee termination (that is, aggregations can only be applied to the results of recursive queries after they have completed, not to results within the bodies of recursive queries).
+The lack of alignment among commercial RDBMS renders a single, unified formal semantics untenable as a foundation for correctness checking of recursive queries; practical systems must instead support a flexible and composable framework that adapts to the specific capabilities of each backend.
+3. How do recursive queries “go wrong”?
+In this section, we classify the problems that arise with recursive queries into three areas based on the emergent database behavior and define the mathematical query properties responsible for each class of error.
+As illustrated in Figure 1, the three behaviors targeted are runtime exception (B1), incorrect results (B2), and nontermination (B3).
+We define six mathematical query properties: range-restriction (P1), monotonicity (P2), mutual-recursion (P3), linearity (P4), set-semantics (P5), and constructor-freedom (P6), and show how each behavior B1-B3 may result from violating one or more of P1-P6, providing an example query (shown in Figure 3) for each query Q1-Q3 in Figure 1.
+Lastly, we discuss the cases where a user may want to selectively apply or relax the restriction of each property, even if the RDBMS supports it.
+3.1. Recursive Query Runtime Exception (B1)
 Bill-of-materials
 Transitive closure
-Transitive closure
+Same-Generation
 Examples of non-monotonic, non-linear, and non-set semantic queries.
+Language-integrated query targets many problems associated with query writing: datatype or schema errors, for example, falsely assuming a table to have a particular column name or data type; syntactic errors like simple typos; security vulnerabilities such as SQL injection; and structural mistakes like HAVING without GROUP BY.
+Without language integration, the RDBMS query compiler will identify these errors and throw an exception, which will be a runtime error for the application.
+These types of errors are already well-covered by existing language-integrated query techniques that are complementary to our approach (Cheney et al., 2013).
+However RDBMS query compilers throw recursion-specific exceptions that are not caught by existing techniques.
+Example.
+The SQL standard specifies that only monotonic queries are supported.
+Practically, this means that all non-monotonic operators, i.e., aggregation or negation, must happen between recursive SQL queries to guarantee termination.
 For example, Figure 3 shows the widely-used Bill-of-Materials problem (Wang et al., 2020) that models items sold by a business that are made out of sub-parts.
-We use the relation SubParts(part, sub) for each item a business sells and its sub-parts (and sub-part sub-parts, etc.); and the relation BasicParts(part, days) for base parts and how many days it takes for that part to arrive from a supplier.
+We use the relation SubParts(part, sub) for each item a business sells and its sub-parts (and sub-sub-parts, etc.); and the relation BasicParts(part, days) for base parts and how many days it takes for that part to arrive from a supplier.
 The query waitFor determines when a part will be ready, given that a part will be ready on the day the last subpart arrives.
-Naively, this query can be represented as Figure 3.
 However, the presence of the MAX aggregation within the body of the recursive query definition means that the query is not monotonic under the ordering of set inclusion.
-The SQL standard specifies that only monotonic queries are supported.
-This constraint is usually checked by the query engine, so the emergent behavior is a run-time error for the application as it must wait for the round-trip time for the query to be sent to the database and the error returned.
-All widely-used commercial databases and Datalog engines officially support only monotonic operations within recursion, so for the query in Figure 3 to pass a database query compiler check, the MAX aggregation must be moved out of the recursive query.
-The set of database query errors that are unrelated to recursion, for example datatype errors, are well covered by previous work in language-integrated query that is complementary to our approach (Cheney et al., 2013).
-Therefore, the mathematical property that leads to recursive query runtime exceptions is non-monotonicity.
+This constraint is usually checked by the query compiler, so the emergent behavior is a runtime error for the application as it must wait for the round-trip time for the query to be sent to the database and the error returned.
+Most widely-used commercial RDBMS and Datalog engines officially support only monotonic operations within recursion, so for the query in Figure 3 to pass a database query compiler check, the MAX aggregation must be moved out of the recursive query.
+Therefore, it is the violation of property P2: monotonicity that leads to behavior B1: recursive query runtime exception on the systems that do not support non-monotonic recursion.
+Queries that violate P1: range-restriction will also always cause runtime exceptions.
+Additionally, many RDBMS do not support mutual or non-linear recursion, so queries that violate property P3: mutual-recursion or P4: linearity will also result in a runtime error for those systems.
+In Figure 1, queries that violate P1, P2, P3, or P4 and are checked by the query compiler are represented by query Q3.
 However, recent advances in recursive query engines have shown that some forms of aggregation (Wang et al., 2020; Abo Khamis et al., 2022) are permissible within recursion without losing termination guarantees.
-While this has not yet been implemented widely in commercial systems, it shows when a user may still want to “turn off” the monotonicity constraint for certain backends.
-2.2.2. Incomplete Results
-The query to compute the transitive closure of a directed graph can be expressed in Datalog with two rules: and .
+While this has not yet been implemented widely in commercial systems, it shows when a user may still want to “turn off” the monotonicity constraint for certain backends without risking nontermination.
+3.2. Incorrect Results (B2)
+A linearly recursive function is a recursive function that contains exactly one recursive call.
+Similarly, the SQL standard defines a linear query to be a query that references the recursive relation once.
+Around the time that SQL’99 was preparing to add recursive queries to SQL, there was a belief that “most ‘real life’ recursive queries are indeed linear” (Bancilhon and Ramakrishnan, 1986).
+While this belief is no longer widely held, this assumption is built into the implementation of many RDBMS.
+Example.
+Figure 3 shows an example of a non-linear query, where the recursive relation path is referenced twice in the body of the query.
+This query is represented by Q2 in Figure 1.
+The reason why RDBMS may return incorrect results is due to an internal optimization that works only for linear queries and is best illustrated with an example: the query to compute the transitive closure of a directed graph can be expressed in Datalog with two rules: and .
 This Datalog program is equivalent to the SQL query in Figure 3.
-However, given a 3-step input graph, i.e.
+Given a 3-step input graph, i.e.
 {(0, 1), (1, 2), (2, 3)} this query returns a result with 5 rows (the input edges plus {(0, 2), (1, 3)}) in PostgreSQL (v15) and DuckDB (v1.1).
 However, when the equivalent query is run in Datalog, the result has 6 rows (the previous results plus (0, 3)).
-Quickly stepping through the graph shows that 3 is reachable from 0, so why is it not in the RDBMS result?
-Unlike Datalog, SQL requires that all WITH RECURSIVE queries be linearly recursive.
-The way this restriction is often enforced is to limit references to the recursive relation to only once within the recursive subquery.
-However this is a purely syntactic restriction, so simple aliasing can easily prevent this check from happening, while other databases do not check query linearity at all.
-Around the time that SQL’99 was preparing to add recursive queries to SQL, there was a belief that “most ‘real life’ recursive queries are indeed linear” (Bancilhon and Ramakrishnan, 1986).
-While this belief is no longer widely held, this assumption is built into the implementation of many database systems.
-Within Datalog engines that use the Semi-Naive evaluation algorithm (Ullman, 1989), for the transitive closure example, at each iteration the “new” data is computed with the subquery expressed in the unnamed relational algebra: where represents the data discovered in the previous iteration, and represents all the data discovered so far over the course of iteration.
+Quickly stepping through the graph shows that 3 is reachable from 0, so the result returned from the RDBMS is incorrect.
+The reason for this is within Datalog engines that use the Semi-Naive evaluation algorithm, for the transitive closure example, at each iteration the “new” data is computed with the subquery expressed in the unnamed relational algebra: where represents the data discovered in the previous iteration, and represents all the data discovered so far.
 However, if a query is guaranteed to be linear, then the query can be optimized to read only , so the previous query becomes which may cause the algorithm to terminate too early for non-linear queries.
-Therefore, the mathematical property that leads to incomplete results is non-linearity.
-As there are modern databases that do not perform any checks that queries are linear and instead allow the query to return incorrect results, it is of utmost importance that users be prevented from unknowingly sending queries that will silently return invalid results.
-However recently, MariaDB claimed to add support for non-linear queries (Shalygina and Novikov, 2017), so a user may wish to turn off this constraint depending on their database backend.
-2.2.3. Nontermination
-WITH RECURSIVE expands the expressive power of SQL to be Turing-complete, so it is possible to represent infinite loops and nontermination.
-Assuming the SQL specification is followed, that is, all queries are monotonic and linear, what causes nontermination?
-Given the query is monotonic over set inclusion, nontermination is the result of infinitely growing intermediate relations.
-Datalog queries are guaranteed to generate only finite relations, given they are correctly range-restricted over a finite domain (Gottlob et al., 1989).
-There are a few key differences in the semantics of RDBMS compared to Datalog engines that cause the mathematical properties that provide termination guarantees to be violated.
-Range-restriction is defined as the constraint that all variables in the head of a Datalog rule must appear in the body of the rule.
-Range-restriction applied to SQL queries requires the project clause of each recursive query to contain only constants or references to columns present in relations in the FROM clause.
-However SQL queries are not limited to a finite domain, as column-level “constructor” operations like addition or string concatenation can arbitrarily introduce new elements.
-The Datalog literature refers to queries that do not contain operations that can infinitely grow the domain as having the constructor-freedom property (Arntzenius and Krishnaswami, 2016).
-Finally, the bottom-up evaluation of logic programs assumes set semantics, which provides termination guarantees at the cost of a performance penalty (checking for and removing duplicates).
-However, RDBMS use bag semantics unless otherwise specified, so a reachability query using bag semantics over data containing cycles will repeatedly re-discover the same paths.
-Only certain operators like DISTINCT or UNION will use set semantics.
-For example, given a linearly recursive transitive closure query, and a cyclic input graph, i.e.
-{(0, 1), (1, 0)}, the query in Figure 3 will infinitely recur.
-Therefore, the mathematical properties that provide termination guarantees are constructor-freedom and set-semantics.
+Terminating early may cause the RDBMS to return only partial results, or when nested within an outer query can lead to fully incorrect results.
+The SQL specification defines behavior only for WITH RECURSIVE queries that are linearly recursive, leaving the behavior of non-linear queries undefined.
+Some RDBMS attempt to reject non-linear queries by limiting references to the recursive relation to only once within the recursive subquery.
+However this is a purely syntactic restriction, so simple aliasing can easily prevent this check from happening, while other databases do not check query linearity at all and allow non-linear queries to execute and silently return incorrect results.
+Systems that allow mutual recursion and implement it via inlining will show similar behavior and return incorrect results.
+Therefore, it is the violation of properties P4: linearity or P3: mutual-recursion that leads to behavior B2: incorrect results on systems that allow non-linear or mutual-recursion.
+As there are modern databases that do not perform checks for P4 and P3, it is of utmost importance that users be prevented from unknowingly sending queries that will silently return invalid results.
+However, recently MariaDB claimed to add support for mutual and non-linear queries (Shalygina and Novikov, 2017), so a user may wish to turn off this constraint depending on their RDBMS.
+3.3. Nontermination (B3)
+WITH RECURSIVE expands the expressive power of SQL to be Turing-complete, so it is possible to express queries that will never terminate.
+Recursive queries can be computationally expensive, and users may be left unsure if their queries will eventually terminate, given enough time, or never terminate.
+Assuming the SQL specification is followed, that is, all queries are monotonic and linear, nontermination is a consequence of infinitely growing intermediate relations.
+Example.
+The bottom-up evaluation assumes set semantics (P5), which prevents duplicate tuples from causing the intermediate relations to grow infinitely.
+Yet RDBMS use bag semantics unless otherwise specified, so a reachability query using bag semantics over data containing cycles will repeatedly re-discover the same paths, leading to infintely growing intermediate relations.
+Only certain operators like DISTINCT or UNION use set semantics.
+For example, the query in Figure 3 finds all descendants of a person (‘A’) that are of the same generation (2) over a table Parents(parent, child).
+Given cyclic input, i.e.
+{(A, B), (B, A)}, this query will infinitely recur.
+Relations can also grow infinitely from non-duplicate tuples if the domain of the query is not finite.
+SQL queries are not limited to a finite domain, as column-level “constructor” operations like addition or string concatenation can arbitrarily introduce new elements.
+The query in Figure 3 contains the + operator, which can produce values not in the input domain and can cause nontermination.
+Therefore, it is the violation of properties P5: set-semantics, or P6: constructor-freedom that leads to behavior B3: nontermination.
+On systems that do not reject non-monotonic operations within recursive queries, property P2: monotonicity can also lead to nontermination.
+Nonterminating queries that violate these properties are represented by Q3 in Figure 1.
+Set-semantic recursive queries are guaranteed to generate only finite relations, given they are correctly range-restricted over a finite domain.
 However, not all queries require set semantics to terminate, although given only the query, it is undecidable whether a query will terminate when using less restrictive duplicate checks (Maher and Ramakrishnan, 1989).
@@
 Other RDBMS require the use of UNION ALL between the base and recursive case of a recursive query, for example SQLServer and OracleDB, and provide alternative language structures to check for infinite recursion (e.g.
-CYCLE in OracleDB and PostgreSQL v14+).
+CYCLE in OracleDB and PostgreSQL v14+ checks if there is a cycle in the query results).
 Users may also wish to turn off the constructor-freedom constraint, as not all constructors lead to infinitely growing domains.
-Range-restriction, however, will always be enforced.
-Given the high penalty of an infinitely recurring query, both on the application and the database itself, as other users may also see reduced performance from cross-query interference, it is important to give users the ability to prevent infinitely recurring queries, even at the cost of possible performance overhead or limits to expressibility.
+Given the high penalty of an infinitely recurring query, both on the application and the RDBMS, as other users may also see reduced performance from cross-query interference, it is important to give users the ability to prevent infinitely recurring queries, even at the cost of possible performance overhead or limits to expressibility.
 4. Overview of Embedded Safe Recursion
+To avoid generating unsafe SQL queries that cause behaviors B1-B3, we must design our system to reject queries that violate properties P1-P6.
+Yet, due to the fractured support for recursion in relational databases, as well as the cases where a user may want to relax certain constraints even for a single RDBMS, for our system to be practically and immediately useful it must be possible to pick-and-choose independently and composably which properties should be constrained at any given time.
+For example, a user of PostgreSQL could choose to constrain mutual-recursion, linearity, and monotonicity but leave the constructor-freedom and set-semantics properties unconstrained.
+Another user should be able to allow bag semantics for queries where they are certain there are no cycles in the data, and set semantics for when there are cycles.
+Constraint-independence informs every aspect of the design and evaluation of our approach.
 RestrictedExpr and RestrictedQuery types and the connection to each query property.
-In Section 2.2, we established the connection between certain problematic database behavior and the mathematical properties of (1) range-restriction, (2) monotonicity, (3) linearity, (4) set-semantics, and (5) constructor-freedom.
-To avoid generating SQL queries that cause this unwanted behavior, we must design our system to reject queries that do not hold these properties independently.
-Figure 4 illustrates how the types representing relations and queries (RestrictedQuery[A, C, D]), and expressions on columns (RestrictedExpr[A, S]) track these properties.
-Each query is built by applying combinators like map and flatMap on relations.
-The RestrictedQuery type represents relations that are recursively defined, the A type parameter represents the column-type of the relation, the C type represents the category, e.g.
-if a query uses bag semantics and can contain duplicates, the D type represents the set of dependency edges between recursive relations.
-The RestrictedExpr type represents expressions that operate over columns of recursive relations, with A representing the result type and S representing the shape of the expression, e.g.
-if it contains any non-monotonic operations.
-Section 4 shows how our system captures these semantic properties of queries and Section 5 uses the host-language type system to translate them into syntactic checks.
-The goal of TyQL is not to provide evaluation semantics for recursive queries because database backends vary widely in their interpretation of recursive queries, but there is one consistency across all evaluation engines: the precedence graph of relation dependencies.
-Therefore, we design our type system to target properties on this graph.
-Intuitively, each of the five properties can be understood as properties on the program graph of the recursive query.
-Relations are represented with nodes, dependencies on source relations are represented as directed edges, and recursive queries are subgraphs with cycles.
-We provide an abstract interpretation that maps terms in to a graph representation in the appendix.
-4. Enforcing Mathematical Properties
-with
+Section 5 describes how the type system captures properties P1-P6 of queries, showing first how each constraint can be implemented independently, then composed into a unified type system.
+Figure 4 shows an overview of the connection between the types that represent queries and expressions on columns and the targeted properties P1-P6.
+The A type parameter represents the column-type of the relation, and is used to enforce P1 (range-restriction).
+For example, a query that returns two columns “age” and “name” will have type (age: Int, name: String).
+The S type parameter represents the shape of the expression, e.g.
+if the expression is monotonic or non-monotonic and is used to enforce P2 (monotonicity).
+The function that constructs RestrictedQuery types, fix, enforces P3 (mutual-recursion).
+The D type parameter represents the set of dependency edges between recursive relations and is used to enforce P4 (linearity).
+The C type parameter represents the category, e.g.
+bag or set semantics, and is used to enforce P5 (set-semantics).
+The RestrictedExpr type represents expressions that are allowed within recursive queries and is used to enforce P6 (constructor-freedom).
+Section 6 describes how to translate each restriction into a syntactic check using the host programming language’s type system.
+The objective of is not to define a full evaluation semantics for recursive queries, which would require either adopting significant simplifying assumptions or excluding features that are essential for practical applicability.
+Instead, addresses the more immediate and impactful problem of reasoning about safety in the context of real-world systems, where safety checks on recursive queries are shaped by the needs of users and inconsistent RDBMS implementations.
+In the appendix, we provide an abstract interpretation that maps terms in to a graph representation that abstracts away differences in RDBMS semantics.
+This approach enables to apply static safety checks that generalize across engines, without relying on a fixed or idealized semantics.
+5. Safe Recursion with
 In this section, we present the calculus for recursive queries.
-We show how the mathematical properties of monotonicity, linearity, set-semantics, range-restriction, and constructor-freedom are encoded into the type system independently so that the restrictions can be composed depending on the targeted database system.
-Because RDBMS have fragmented support for recursion, yet offer significant performance benefits, it is essential that the type system allow restrictions to be tuned and not overly constrict real-world queries.
+We show how the query properties P1-P6 are encoded into the type system independently so that the restrictions can be composed depending on the targeted database system.
 We first present the DSL type system in Section 5 and then show how to extend it to a 2-level DSL and host language type system in Section 6.
 5.1. Syntax and Base Type System
+The base of extends T-LINQ (Cheney et al., 2013).
+T-LINQ, as well as the Nested Relational Calculus (NRC) (Buneman et al., 1995), established the structural equivalence between SQL-style SELECT-FROM-WHERE and combinator-style flatMap, filter, map, etc.
+expressions.
+The recursive query syntax is built on previous work on the NRC with a bounded fixpoint operator (Suciu, 1997).
+BOM query
+To illustrate an example, Figure 5 shows a recursive query on the BOM database from Section 3 in SQL and .
+The appendix also includes the queries of Figure 3 in .
+In TyQL, for-comprehensions desugar to combinators but we do not include for-comprehensions in to keep the syntax minimal.
+Recursive queries are composed of base cases that contain no references to recursive relations, and recursive cases where references to the recursive relations are allowed.
+Although some databases do not syntactically require a base case, we enforce this in as recursion without a base case leads to undefined behavior.
 Syntax of Types and Terms
-BOM query
-The non-recursive part of extends T-LINQ (Cheney et al., 2013).
 The syntax of types and terms and the typing judgments are presented in Figure 6.
 ranges over variables.
-The column types represent database column types, Named-Tuples represent row types, represents a query that returns a collection of rows of type A, and represents an aggregation that returns a single scalar result of type A.
+Database column types are represented with the column types, multi-column row types with Named-Tuples, represents a relation or query that returns a collection of rows of type A, and represents an aggregation that returns a single scalar result of type A.
 Tuples and function types are used to construct the combinators () that represent database operations.
-The combinators follow the precedent set by T-LINQ, and we add to represent database-level recursion.
-We refer to “column-level expressions” as any expression that would go into the filter or project subclause of a SQL query: for map, it would be SELECT a + 1 FROM table vs.
-for aggregate it would be SELECT sum(a) + 1 FROM TABLE.
+The combinators follow the precedent set by T-LINQ, and we add fix to represent database recursion.
+“Column-level expressions” refer to any expression that can go into the filter or project clause of a query: for map, it would be SELECT a + 1 FROM R vs.
+for aggregate it would be SELECT sum(a) + 1 FROM R.
 Both a + 1 and sum(a) + 1 are column-level expressions.
-Query-level expressions refer to the entire query expression and operations that take relations, e.g.
+“Query-level expressions” refer to the full query expression and operations that combine subqueries, e.g.
 union.
+The DSL syntax does not include application of functions defined by the user (i.e., ).
+Function types can be passed to the constructs in but do not get executed until a later stage.
+DSL Typing rules with only range-restricted recursion
 Following the convention set by T-LINQ, we assume a signature that maps each constant c and operator to the typing rule.
@@
 union).
+In Figure 6 and the typing rules we use the syntax as a stand-in for all operations, even though some operations like or use infix syntax.
 Examples of operations stored in are presented in Figure 6.
-DSL Typing rules with only range-restricted recursion
-We build our syntax on the Nested Relational Algebra extended with a “bounded fixpoint” operator (Suciu, 1997).
-To illustrate an example, in Figure 2 we show a recursive query on the BOM database from Section 2.2 in SQL and .
-In the TyQL implementation, for-comprehensions desugar to combinators but we do not include for-comprehensions in to keep the syntax minimal.
-In SQL, a recursive query is composed of a “base case” that contains no references to recursive relations, and a “recursive case” where references to the recursive relation are allowed.
-Although some databases do not syntactically require a base case, we enforce this in because recursion without a base case leads to undefined behavior.
 Figure 7 shows the typing rules for the DSL, which closely follows the precedent type system in T-LINQ, with the addition of fix.
@@
 The fix function defines recursive relations within a single stratum.
-For , fix takes as arguments a tuple of “base case” queries and a function .
-The function takes as arguments a tuple of references to the recursive relations being defined and returns a tuple of “recursive case” definitions .
+For , fix takes as arguments a tuple of base case queries and a function .
+The function takes as arguments a tuple of references to the recursive relations being defined and returns a tuple of recursive case definitions .
 Each is composed from the recursive references (and any relations in scope that are defined outside the body of fix) using the combinators in .
-In the next section, we show how we apply the type system to independently identify and prevent the mathematical properties identified in Section 2.2.
-4.2. Range-restricted Recursion
-Range-restriction is the constraint that every variable appearing in the head of a rule appears in the body of that rule, to avoid rules that yield infinite relations.
-Because this constraint does not depend on the semantics of the database system, it is encoded into the basic FIX-D typing rule in Figure 7 by enforcing the recursive references and the recursive definitions all have the same type, and that is a Named-Tuple.
+In the next section, we show how we apply the type system to independently identify and prevent properties P1-P6.
+We start with P1 and include it in the base type system as there are no cases where a user would want to turn off range-restriction.
+5.2. P1: Range-restricted Recursion
+Range-restriction is the constraint that every variable appearing in the left-hand-side of a Datalog rule appears in the right-hand-side, to avoid rules that yield infinite relations.
+Because this constraint does not depend on the semantics of the database system, it is encoded into the basic FIX-D typing rule in Figure 7 by enforcing the recursive references and the recursive definitions all have the same type, , and that is a Named-Tuple.
 As Named-Tuples must have the same key and value types, order, and arity to be considered the same type, this restriction enforces that all variables in the head of a the rule (e.g.
@@
 Definition 1.
-A function such that holds the range-restricted property if .
-4.3. Monotone Recursion
+A function such that holds the range-restricted property P1 if .
+5.3. P2: Monotone Recursion
+Operations that are non-monotonic under set inclusion are defined as scalar operations.
+For example, the query SELECT max(a) + 1 FROM R is non-monotonic and produces a single value, while SELECT a + 1 FROM R is monotonic and nonscalar and produces one value per row in R.
+Both scalar and nonscalar expressions can be constructed from each other: + is nonscalar while sum is scalar but the expressions sum(a + 1) or sum(a) + 1 are both valid and should produce an expression of type scalar, while SELECT x + (SELECT max(y) FROM R1) FROM R2 should produce an expression of type nonscalar (as the subquery does not change that the full expression is monotonic).
+The type system must be able to distinguish the shape of the entire expression (no matter how nested) so that well-formed terms of type Query are guaranteed to return nonscalar results and hold P2.
 | Syntax | |||||
-Operations that are non-monotonic under set inclusion are defined as scalar operations.
-For example, the query SELECT max(a) + 1 FROM TABLE is non-monotonic and produces a single value, while SELECT a + 1 FROM table is monotonic and nonscalar.
-Both scalar and nonscalar expressions can be constructed from each other: + is nonscalar while sum is scalar but the expressions sum(a + 1) or sum(a) + 1 are both valid and should produce an expression of type scalar, while SELECT x + (SELECT max(y) FROM table1) FROM table2 should produce an expression of type nonscalar (as the subquery does not change that the full expression is monotonic).
-The type system must be able to distinguish the shape of the entire expression (no matter how nested) so that well-formed terms of type Query are guaranteed to return nonscalar results and are monotonic under set inclusion.
 The required changes to the type system to restrict recursive queries to monotone operations are shown in Figure 8.
@@
 By EXPR-OP-D and AGG-OP-D, any expression that contains an arbitrarily nested sub-expression with type parameter will be of type (via Shape).
-For example:
-The expression on the left contains a nested scalar operation (sum) so the resulting type will be , while the expression on the right contains only nonscalar expressions, so even with the same operator () the result type is The MAP-D and FLATMAP-D rules are refined to only accept functions that contain non-scalar expressions, while AGGREGATE-D and GROUPBY-D only accept functions that contain at least one scalar sub-expression.
-With these restrictions, expressions of type Query in are guaranteed to contain only monotonic operations.
-The monotonicity property is relevant only between the relations in a single stratum (Bancilhon and Ramakrishnan, 1986).
-To limit the monotonicity restriction to only recursive references, we introduce a RestrictedQuery type and refine FIX-D so that the arguments to must be of type RestrictedQuery.
-Crucially, RestrictedQuery can only be derived by calling combinators on the arguments passed to the function given to fix as it is not available in the syntax.
+For example, depending on the subexpression, the EXPR-OP-D rule applied to the ”+” operator can produce an expression of type Scalar or NonScalar:
+The expression on the left contains a nested scalar operation (sum) so the resulting type will be , while the expression on the right contains only nonscalar expressions, so even with the same operator () the result type is .
+The MAP-D and FLATMAP-D rules are refined to only accept functions that return non-scalar expressions, while AGGREGATE-D and GROUPBY-D only accept functions that return at least one scalar sub-expression.
+With these restrictions, expressions of type Query are guaranteed to contain only monotonic operations, while only expressions of type Aggregation may contain non-monotonic operations.
+As discussed in Section 2, the monotonicity property only needs to restrict aggregation between the relations in a single stratum, i.e., within recursive queries to recursive references.
+To limit the monotonicity restriction to only recursive references, we introduce a RestrictedQuery type and refine FIX-D so that the arguments and return type of must be of type RestrictedQuery.
+Crucially, RestrictedQuery can only be derived by calling combinators on the arguments passed to the function given to fix as it is not available in the syntax and therefore can only be in scope within the body of fix.
 We update the MAP-D, FLATMAP-D, FILTER-D, and REL-OP-D so if they are passed any arguments of type RestrictedQuery the result type will be RestrictedQuery.
 To reduce the number of rules in Figure 8, we define a meta-syntax helper to abstract away the differences between rules for combinators that take RestrictedQuery and Query.
-Without , we could equivalently define 4 rules for flatMap: if and , the result type will be , however if either or are RestrictedQuery the result will be .
-As AGGREGATE-D and GROUPBY-D are valid only on Query types, will reject non-monotonic operations on recursive references.
+Without , we could equivalently define 4 rules for flatMap: if and , the result type will be , however if either or the return type of are RestrictedQuery the result will be .
+As AGGREGATE-D and GROUPBY-D are valid only on Query types, aggregations cannot be applied to recursive references and will reject non-monotonic operations on recursive relations within recursive queries.
 Definition 2.
 A function such that holds the monotonicity property if .
-4.4. Linear Recursion
+5.4. P3: Mutually-recursive Recursion
+restricts mutually-recursive queries by simply limiting the number of arguments to the fix function to a single argument.
+Single-direction dependencies between recursive relations can be defined using multiple calls to fix.
+Definition 3.
+A function such that holds the mutually-recursive property P3 if .
+5.5. P4: Linear Recursion
 fix restricted to be linear
-We approach the problem of encoding linear recursion in the type system with a linear type with respect to recursive references.
-To accomplish this, we split the problem into two sub-constraints: an affine check (every relation is never used more than once) and a relevant check (all relations are used in at least one recursive definition).
-The affine check is per-relation, while the relevant check is for all relations, as not every relation needs to use every other relation within a single stratum.
+We approach the problem of encoding linear recursion in the type system with a linear type with respect to recursive relations.
+To accomplish this, we split the problem into two sub-constraints: an affine check (every recursive relation is never used more than once to define a recursive relation) and a relevant check (all recursive relations are used in at least one recursive definition).
+The affine check is per-relation, while the relevant check is for all relations defined within a single fix call, as not every relation needs to use every other relation within a single stratum.
 The extensions to the type system to restrict queries to linear recursion are shown in Figure 9.
 A type parameter of type is added to to model the dependencies of each query.
+We use to indicate multiset union, so duplicates are maintained, and equivalence using does not take order into account.
 Uniqueness of references is enforced by using the argument position within the fix function: for a recursive function , will be a tuple containing the constant integer type , and will contain a tuple with the constant integer type .
@@
 This ensures that if fix is called within the scope of references from another fix function, the inner fix cannot return terms derived from the outer function’s parameters.
+Multi-relation operations like flatMap or union collate recursive references in the result relation.
 For example:
 For the affine check, the dependencies of each recursive definition must contain no duplicates.
+This check is implemented by taking the length (indicated with ) of and requiring that is the same as the length of the union of : .
 For the relevant check, all parameters must appear at least once in all .
-This is implemented by using the set of constant integer types from 1 to the number of arguments to fix decorated with to check that all elements are present at least once in all the s () and never more than once in each ().
+This is implemented by using the set of constant integer types from 1 to the number of arguments to fix decorated with to check that all elements are present at least once in all the s: .
 The REL-OP-D rule is updated so if any arguments are RecursiveQuery, the result will be a RecursiveQuery that collates (only) the dependencies of the recursive relations.
@@
 Definition 4.
-If then holds the linearity property if it holds both the relevant and affine properties. The relevant property holds if and the affine property if .
-4.5. Set semantic Recursion
-| Syntax | |||||
+If then holds the linearity property P4 if it holds both the relevant and affine properties. The relevant property holds if and the affine property if .
+5.6. P5: Set-semantic Recursion
+| Syntax | ||||||||
 fix restricted to be set semantic
-To restrict fix to allow only set semantics within recursive queries, we track the semantics of each operation in a type parameter, similarly to how we tracked monotonicity.
-We update so that all relation-level operations reflect their semantics.
+To restrict fix to allow only set semantics within recursive queries, we track the semantics of each query-level operation in a type parameter, similarly to how we tracked monotonicity in column-level operations.
+We update so that all query-level operations reflect their semantics.
 The required changes to the syntax and typing rules are presented in Figure 10.
@@
 Definition 5.
-A function such that holds the set-semantic property if .
-4.6. Constructor-Freedom
+A function such that holds the set-semantic property P5 if .
+5.7. P6: Constructor-Freedom
 Constructors, e.g.
@@
 Definition 6.
-A function such that holds the constructor-freedom property if and is of type RestrictedExpr.
+A function such that holds the constructor-freedom property P6 if and is of type RestrictedExpr.
 5.8. Nesting in
 The problem of deeply nested datatypes and query avalanches is well covered by previous work in NRC and T-LINQ.
-The unnesting and normalization approach taken by T-LINQ is applicable to non-recursive queries, or solely within the bodies of recursive queries, but it is not applicable between non-recursive and recursive subqueries because the fixpoint introduces a strict evaluation boundary—often referred to as a ‘pipeline breaker’, which prevents the same normalization approach.
+The unnesting and normalization approach taken by T-LINQ is applicable to non-recursive queries, or solely within the bodies of recursive queries, but it is not applicable between non-recursive and recursive subqueries because the fixpoint introduces a strict evaluation boundary - often referred to as a ‘pipeline breaker’, which prevents the same normalization approach.
 Therefore, we have chosen to simplify the syntax of to restrict nested column types, although in our implementation (discussed in Section 7) nested datatypes are supported.
@@
 We reuse the syntax of types and terms in Figure 6 for the host language, with a few key additions.
+Lambda application, i.e., , is added to the host-language syntax with the typing rule APP so users can define and use arbitrary functions.
 DSL types are represented in the host language by wrapping them in Expr types, so if a DSL expression represents a database row of type then the type of this expression in the host language will be .
@@
 The typing rules for in Figure 11 combine all the restrictions presented in Figures 8, 9, 10 with the addition of rules to handle interactions between the DSL and host language.
-The typing rule for the fully-restricted host language fix is shown in Figure LABEL:fig:fix-impl.
+The typing rule for the fully-restricted host language fix is shown in Figure LABEL:fig:fix-impl (to restrict mutual recursion, we limit the meta-syntax to ).
 The DSL typing judgments must also be updated slightly.
@@
 In this section, we describe the implementation of TyQL, our type-safe embedded query library based on .
-Figure LABEL:fig:fix-impl shows the implementation of fix in TyQL with all restrictions applied and the typing rule from with all restrictions applied, renamed to match the variable names in TyQL for readability.
+Figure LABEL:fig:fix-impl shows the implementation of fix in TyQL with all restrictions applied except mutual recursion and the typing rule from , renamed to match the variable names in TyQL for readability.
 We include the side-by-side comparison to emphasize that our implementation and our theory are closely knit.
@@
 This method is used as an implicit conversion between Named-Tuple-of-Expr and Project[A] (line 8).
-TyQL reifies the constant string keys in Named-Tuples using type tags (ResultTag) so that the generated queries are more readable, for example t.map(r => (maxCount = max(r.count)) generates the SQL query SELECT MAX(count) as maxCount FROM t.
+TyQL reifies the constant string keys in Named-Tuples using type tags (ResultTag) so that the generated queries are more readable, for example r.map(r => (maxCount = max(r.count)) generates the SQL query SELECT MAX(count) as maxCount FROM R.
 Hierarchy of types in TyQL.
@@
 TyQL uses the annotation @implicitNotFound to customize error messages, so that the user can see which constraint failed: i.e., if their recursive query was affine but not relevant.
-For systems that do not support mutual recursion, we do not have to include this constraint in the type system because we can simply limit the number of arguments to the fix function to a single relation, which prevents the construction of mutually recursive queries.
+To restrict mutual recursion in Figure LABEL:fig:fix-impl, we could tweak fix such that QT, DT, and RQT must be subtypes of Tuple1 (tuples of length 1).
+However, we prefer to provide a more erognomic implementation where fix is a method of Query, so that users can simply write edges.fix(e -> ...).
 Unique query references are implemented using constant integer types and unique types are implemented using anonymous classes and value types.
-Figure LABEL:fig:fix-impl shows the fully restricted fix, but the restrictions can be made configurable: for example, if a user wants a fix function with all restrictions except set-semantics, they simply call val bagFixFn = fix((ConstructorFree(), Monotone(), BagResult(), Linear())).
+The restrictions of P1-P6 can be made configurable: for example, if a user wants a fix function with all restrictions except set-semantics, they call val bagFixFn = fix((ConstructorFree(), Monotone(), BagResult(), Linear())).
 This is implemented by adding additional type parameters for each restriction [RCF <: ConstructorFreedom, RM <: MonotoneRestriction, RC <: ResultCategory, RL <: LinearRestriction] to line 2 and an additional argument to line 3: (options: (RCF, RM, RC, RL)).
@@
 A similar approach is applied to sequences of flatMap, for example the query t1.flatMap(r1 => t2.flatMap(r2 => t3.map(r3 => (r1 = r1.x, r2 = r2.x, r3 = r3.x)))) is flattened into a single JOIN node (as opposed to a binary, left-deep tree).
+With respect to query construction, TyQL doesn’t introduce any additional overheads from JVM boxing beyond what Scala does when working with generic ADTs or case classes.
+With respect to results sent back from the database, all boxing is handled by the JDBC driver regardless of if the query was constructed with strings or with TyQL.
 8. Evaluation
-| ⬇ // TyQL: val base = Table[(dst:Int, cst:Int)]("base") val edges = Table[(src:Int, dst:Int, cst:Int)]("edge") base.fix(sp => edges.flatMap(edge => sp.filter(s => s.dst == edge.src) .map(s => (dst = edge.dst, cst = s.cst + edge.cst)) .distinct) .aggregate(s => (dst = s.dst, cst = min(s.cst))) .groupBy(s => (dst = s._1.dst)) | ⬇ -- SQL: CREATE TABLE edge (src INT, dst INT, cst INT); CREATE TABLE base (dst INT, cst INT); WITH RECURSIVE path AS ( (SELECT * FROM base) UNION (SELECT edge.dst, path.cst + edge.cst FROM edge, path WHERE path.dst = edge.src)) SELECT path.dst, MIN(path.cst) FROM path GROUP BY path.dst |
-SSSP Query.
-| Query | Description | Linear | Mono-tone | Set | Mut-ual | CF |
-| Even-Odd | ◼ Mutual recurive program generates even/odd numbers. | ✓ | ✓ | ✓ | ✓ | ✗ |
-| CSPA | ⚫ Graspan’s Context Sensitive Pointer Analysis, static analysis for identifying aliases. (Wang et al., 2017) | ✗ | ✓ | ✓ | ✓ | ✓ |
-| Company Control (CC) | ▲ Calculates the complex controlling relationships between companies. (Mumick et al., 1990) | ✓ | ✗ | ✓ | ✓ | ✗ |
-| PointsToCount (PTC) | ⚫ Find the count of objects that a variable of a given name may point to (Abeysinghe et al., 2024) (Java variant). | ✗ | • | ✓ | ✓ | ✓ |
-| Chain Of Trust (COT) | ◼ Security query where two entities trust each other if they are direct friends or if their friends trust each other. | ✓ | ✓ | ✗ | ✓ | ✓ |
-| Java Points To (JPT) | ⚫ Field-sensitive subset-based points to analysis for an object-oriented language. (Madsen et al., 2016) | ✗ | ✓ | ✗ | ✓ | ✓ |
-| Party | ★ Mutually recursive social media graph algorithm.(Wang et al., 2020) | ✓ | ✗ | ✗ | ✓ | ✗ |
-| CBA | ◼ Constraint-based analysis query. (souffle, 2019) | ✗ | • | ✗ | ✓ | ✓ |
-| Single Source Shortest Path (SSSP) | ★ Computes the shortest path from a given source node to all other nodes in a weighted graph. (Wu et al., 2022) | ✓ | • | ✓ | ✗ | ✗ |
-| Ancestry | ▲ Find descendants of a person who are of the same generation. (Bancilhon and Ramakrishnan, 1986) | ✓ | ✓ | ✓ | ✗ | ✗ |
-| Andersen’s Points To (APT) | ⚫ Context-insensitive, flow-insensitive interprocedural pointer analysis. (Fan et al., 2019a) | ✗ | ✓ | ✓ | ✗ | ✓ |
-| All Pairs Shortest Path (APSP) | ★ Compute the shortest paths between all pairs of nodes in a weighted graph. (Wu et al., 2022) | ✗ | ✗ | ✓ | ✗ | ✗ |
-| Graphalytics (TC) | ★ Directed cyclic graph reachability query. Uses list data structure to check for cycles. (Angles et al., 2020) | ✓ | ✓ | ✗ | ✗ | ✗ |
-| Bill of Materials (BOM) | ▲ Business query for how many days it would take to deliver a product made of subparts. (ibm, 2021) | ✓ | • | ✗ | ✗ | ✓ |
-| Orbits | ◼ Orbits of cosmological objects. (souffle, 2019) | ✗ | • | ✗ | ✗ | ✓ |
-| Data Flow | ◼ Models control flow through read/write/jump instructions. (souffle, 2019) | ✗ | ✓ | ✗ | ✗ | ✓ |
-Linear, Set, Mutual, CF (constructor-free): ✓ query has the property.
-Monotone: ✓ no aggregation • stratified aggregation ✗ unstratified aggregation.
 In this section, we evaluate TyQL with respect to real-world use cases and how effectively TyQL can constrain recursive queries without compromising on usability.
 To illustrate the range of real-world queries, we introduce a recursive query benchmark (RQB) of 16 queries taken from open source code repositories and recent publications from various query domains, including business management, program analysis, graph queries, and other classic fixed-point problems.
-We give a survey of modern relational database systems using our benchmark, identifying the ways each query can go wrong and which RDBMS supports each query property or combination of properties.
+We give a survey of modern relational database systems using our benchmark, identifying the ways each query can go wrong and which RDBMS supports each query property or combination of P1-P6.
 We use the RQB to evaluate TyQL with respect to query coverage: how many queries that would go wrong are prevented from compiling, and impact: if TyQL were to allow the query to compile, how it would fail on the database.
@@
 8.1. Recursive Query Benchmark and Survey of RDBMS
-In this section, we present a recursive query benchmark (RQB) comprising 16 representative queries across diverse domains such as business management, program analysis, graph queries, and classic fixed-point problems and explore how each query behaves on different RDBMS.
+| Query | Description | P2 | P3 | P4 | P5 | P6 |
+| Even-Odd | ◼ Mutually recursive program generates even/odd numbers. | ✓ | ✓ | ✓ | ✓ | ✗ |
+| CSPA | ⚫ Graspan’s Context Sensitive Pointer Analysis, static analysis for identifying aliases. (Wang et al., 2017) | ✓ | ✓ | ✗ | ✓ | ✓ |
+| Company Control (CC) | ▲ Calculates the complex controlling relationships between companies. (Mumick et al., 1990) | ✗ | ✓ | ✓ | ✓ | ✗ |
+| PointsToCount (PTC) | ⚫ Find the count of objects that a variable of a given name may point to (Abeysinghe et al., 2024) (Java variant). | • | ✓ | ✗ | ✓ | ✓ |
+| Chain Of Trust (COT) | ◼ Security query where two entities trust each other if they are direct friends or if their friends trust each other. | ✓ | ✓ | ✓ | ✗ | ✓ |
+| Java Points To (JPT) | ⚫ Field-sensitive subset-based points to analysis for an object-oriented language. (Madsen et al., 2016) | ✓ | ✓ | ✗ | ✗ | ✓ |
+| Party | ★ Mutually recursive social media graph algorithm.(Wang et al., 2020) | ✗ | ✓ | ✓ | ✗ | ✗ |
+| CBA | ◼ Constraint-based analysis query. (souffle, 2019) | • | ✓ | ✗ | ✗ | ✓ |
+| Single Source Shortest Path (SSSP) | ★ Computes the shortest path from a given source node to all other nodes in a weighted graph. (Wu et al., 2022) | • | ✗ | ✓ | ✓ | ✗ |
+| Ancestry | ▲ Find descendants of a person who are of the same generation. (Bancilhon and Ramakrishnan, 1986) | ✓ | ✗ | ✓ | ✓ | ✗ |
+| Andersen’s Points To (APT) | ⚫ Context-insensitive, flow-insensitive interprocedural pointer analysis. (Fan et al., 2019) | ✓ | ✗ | ✗ | ✓ | ✓ |
+| All Pairs Shortest Path (APSP) | ★ Compute the shortest paths between all pairs of nodes in a weighted graph. (Wu et al., 2022) | ✗ | ✗ | ✗ | ✓ | ✗ |
+| Graphalytics (TC) | ★ Directed cyclic graph reachability query. Uses list data structure to check for cycles. (Angles et al., 2020) | ✓ | ✗ | ✓ | ✗ | ✗ |
+| Bill of Materials (BOM) [Stratified] | ▲ Business query for days to deliver a product made of subparts with stratified aggregation. (ibm, 2021) | • | ✗ | ✓ | ✗ | ✓ |
+| Orbits | ◼ Orbits of cosmological objects. (souffle, 2019) | • | ✗ | ✗ | ✗ | ✓ |
+| Data Flow | ◼ Models control flow through read/write/jump instructions. (souffle, 2019) | ✓ | ✗ | ✗ | ✗ | ✓ |
+Benchmark domain: ◼ Datalog ⚫ Program analysis ▲ Recursive SQL ★ Graph.
+P2 (monotone): ✓ no aggregation • stratified aggregation ✗ unstratified aggregation.
+P3 (mutually-recursive), P4 (linear), P5 (set-semantic), P6 (constructor-free): ✓ query has the property.
+In this section, we present a recursive query benchmark (RQB) comprising 16 queries across diverse domains such as business management, program analysis, graph queries, and classic fixed-point problems and show how each query behaves on different RDBMS.
 The goal of our benchmark is to simulate a broad range of recursive queries.
-We have identified five discrete properties of recursive queries that have the greatest impact on behavior: (1) linearity, (2) monotonicity, (3) set semantics, (4) mutual recursion, and (5) constructor-freedom.
-All queries are range-restricted.
-We have selected 16 queries to represent classes of queries for each combination of properties 1-4.
-Since constructor-freedom applies independently to any query, we included a mix of queries with this property.
+We have selected 16 queries to represent classes of queries for each combination of properties (except P1 as all queries are range-restricted).
+We excluded trivially safe queries (those violating none of P1–P6) because they pose no risk of causing behaviors B1-B3.
 Identifying the frequency of each query across real-world applications remains future work.
-All queries run and terminate on at least one of the evaluated database systems.
+All queries run and terminate on at least one of the evaluated RDBMS.
 For the monotonicity property, we have selected a mix of queries without aggregation or negation, with stratified-aggregation, and with unstratified aggregation.
@@
 As an example, we illustrate the Single Source Shortest Path query expressed in TyQL and in SQL in Figure 15.
-The full set of queries in the recursive benchmark expressed in both TyQL and the generated SQL are included in the appendix.
-| DuckDB | Postgres | SQLite | MariaDB | What restriction is violated? | |
-| Even-Odd | ✓ | ✗ | ✗ | ◗ | mutual, cf |
-| CSPA | ◗ | ✗ | ✗ | ✗ | linear, mutual |
-| CC | ◗ | ✗ | ✗ | ✗ | monotone, mutual, cf |
-| PTC | ◗ | ✗ | ✗ | ✗ | linear, mutual |
-| COT | ✓ ❘ ▼✳ | ✗ | ✗ | ✗ | set, mutual |
-| JPT | ◗ ❘ ▼✳ | ✗ | ✗ | ✗ | linear, set, mutual |
-| Party | ◗ | ✗ | ✗ | ▼ | monotone, set, mutual, cf |
-| CBA | ◗ ❘ ▼✳ | ✗ | ✗ | ✗ | linear, set, mutual |
-| SSSP | ✓ ❘ ▼✳ | ✓ ❘ ▼ | ✓ ❘ ▼ | ✓ | cf |
-| Ancestry | ✓ | ✓ | ✓ | ✓ | cf |
-| APT | ◗ | ◗ | ✗ | ◗ | linear |
-| APSP | ◗ | ◗ | ✗ | ✓ | linear, monotone, cf |
-| TC | ✓ | ✓ | ✓ | ✓ | set, cf |
-| BOM | ✓ ❘ ▼✳ | ✓ ❘ ▼✳ | ✓ ❘ ▼✳ | ✓ | set |
-| Orbits | ◗ ❘ ▼✳ | ◗ ❘ ▼✳ | ✗ | ▼ | linear, set |
-| Data Flow | ◗ ❘ ▼✳ | ◗ ❘ ▼✳ | ✗ | ▼ | linear, set |
-As shown in Section 2.2, support for recursive queries varies widely across RDBMS.
+The full set of queries in the recursive benchmark expressed in both TyQL and SQL are included in the appendix.
+| ⬇ // TyQL: val base = Table[(dst:Int, cst:Int)]("base") val edges = Table[(src:Int, dst:Int, cst:Int)]("edge") base.fix(path => edges.flatMap(edge => path.filter(s => s.dst edge.src) .map(s => (dst = edge.dst, cst = s.cst + edge.cst)) .distinct) .aggregate(s => (dst = s.dst, cst = min(s.cst))) .groupBy(s => (dst = s._1.dst)) | ⬇ -- SQL: CREATE TABLE Edges (src INT, dst INT, cst INT); CREATE TABLE Base (dst INT, cst INT); WITH RECURSIVE Path AS ( (SELECT * FROM Base) UNION (SELECT Edges.dst, Path.cst + Edges.cst FROM Edges, Path WHERE Path.dst = Edges.src)) SELECT Path.dst, MIN(Path.cst) FROM Path GROUP BY Path.dst |
+SSSP Query.
+| Benchmark | Database Behavior | ||||
+| Query | Violated Properties | DuckDB | Postgres | SQLite | MariaDB |
+| Even-Odd | P3, P6 | ✓ | ✗ | ✗ | ◗ |
+| CSPA | P3, P4 | ◗ | ✗ | ✗ | ✗ |
+| CC | P2, P3, P6 | ◗ | ✗ | ✗ | ✗ |
+| PTC | P3, P4 | ◗ | ✗ | ✗ | ✗ |
+| COT | P3, P5 | ✓ ❘ ▼✳ | ✗ | ✗ | ✗ |
+| JPT | P3, P4, P5 | ◗ ❘ ▼✳ | ✗ | ✗ | ✗ |
+| Party | P2, P3, P5, P6 | ◗ | ✗ | ✗ | ▼ |
+| CBA | P3, P4, P5 | ◗ ❘ ▼✳ | ✗ | ✗ | ✗ |
+| SSSP | P6 | ✓ ❘ ▼✳ | ✓ ❘ ▼ | ✓ ❘ ▼ | ✓ |
+| Ancestry | P6 | ✓ | ✓ | ✓ | ✓ |
+| APT | P4 | ◗ | ◗ | ✗ | ◗ |
+| APSP | P2, P4, P6 | ◗ | ◗ | ✗ | ✓ |
+| TC | P5, P6 | ✓ | ✓ | ✓ | ✓ |
+| BOM | P5 | ✓ ❘ ▼✳ | ✓ ❘ ▼✳ | ✓ ❘ ▼✳ | ✓ |
+| Orbits | P4, P5 | ◗ ❘ ▼✳ | ◗ ❘ ▼✳ | ✗ | ▼ |
+| Data Flow | P4, P5 | ◗ ❘ ▼✳ | ◗ ❘ ▼✳ | ✗ | ▼ |
+As shown in Section 2, support for recursive queries varies widely across RDBMS.
 To get a sense of what classes of queries each system supports, we ran the 16 queries with cyclic input data and with acyclic input data.
 The results of the RQB on four RDBMS are presented in Table 3.
-We represent queries that terminated with the full result with ✓, database error with ✗, incomplete results with ◗, and nontermination with ▼.
+We represent queries that terminated with the full result with ✓, database error (B1) with ✗, incomplete results (B2) with ◗, and nontermination (B3) with ▼.
 If the query exhibited different behavior based on if the input data was cyclic or not, for example if the query terminates correctly with acylic data but not cyclic data, then the query is represented as ✓ ❘ ▼ with the acyclic data on the left and cyclic data on the right.
-We consider “incomplete” results to be data that is missing results that would have been present using a non-SQL version of each algorithm, either via general-purpose imperative programs or Datalog programs.
+We consider “incomplete” results to be data that is missing results present using a non-SQL version of each algorithm, either via imperative programs or Datalog programs.
 The • and ▼ classification indicates that it is possible to find an input dataset that returns incomplete results or does not terminate.
@@
 In this section, we use the RQB to evaluate the effectiveness of TyQL in achieving comprehensive query coverage while preventing queries that may go wrong from compiling, illustrated in Table 3.
-The Ancestry query exemplifies the class of queries that exhibit only the properties fully supported by the SQL specification: it is linear, monotonic, set-semantic and does not contain any mutual recursion.
-Only with constructor-freedom enforced will TyQL reject this query.
-We provide this option to show the completeness of our approach, although we expect most real-world queries to fall into this category and do not anticipate that it will be widely applied because it is very restrictive for real-world use cases.
+The goal of TyQL is to target unwanted database behaviors B1-B3, while the mechanism is via deriving the query properties P1-P6 using the type system.
+P1-P6 are ground truths for each query, regardless if the query is expressed in raw SQL, TyQL, Datalog, or another language.
+With respect to P1-P6, there are no false positives or false negatives because the typing rules presented in Figure 11 are derived directly from Definitions 1- 6.
+The behaviors B1-B3, however, are a property of the semantics of the database backend.
+Therefore, it is only possible to classify queries as positive or negative with respect to a database and TyQL configuration, e.g., what set of properties P1-P6 are enforced by TyQL.
+Table 3 shows how each query can be considered a false positive with respect to a database (represented ✓, e.g.
+no problems during execution), or a true positive (and the impact on the database, represented by B1 ✗, B2 ◗, and B3 ▼), for a set of constraints (the “Violated Properties” column).
+True negatives are the queries that run without problems for the set of properties not violated.
+For example, using Postgres with properties P2, P3, P4 over cyclic data, the TC query is a false positive while the SSSP query is a true positive.
+The Ancestry query exemplifies the class of queries that exhibit only the properties fully supported by the SQL specification: it is linear, monotonic, set-semantic, and does not contain any mutual recursion.
+Only with constructor-freedom (P6) enforced will TyQL reject this query.
+We provide this configuration to show the completeness of our approach, although we expect most real-world queries to fall into this category and do not anticipate that P6 will be widely applied because it is very restrictive for real-world use cases.
 For the queries that return incorrect results, all are either mutually recursive or non-linear.
@@
 Nonterminating queries can have a significant performance impact, both on the application and on the other users of the database, due to interference.
-Some databases allow you to set a max recursion depth to avoid infinite recursion, although it is not obvious how deep to set this without trial and error.
+Some RDBMS allow you to set a max recursion depth to avoid infinite recursion, although it is not obvious how deep to set this without trial and error.
 It is clear that the impact of incorrect results is more damaging than a database throwing an error.
 However, if nontermination is more impactful than incomplete results depends on the context of the application and database system.
-So far, the only relational data management system we have seen to officially include non-linear or mutual recursion is MariaDB (Shalygina and Novikov, 2017), although in our experiments we found that it often returned slightly different results from Datalog engines.
+So far, the only RDBMS we have seen to officially include non-linear or mutual recursion is MariaDB (Shalygina and Novikov, 2017), although in our experiments we found that it often returned slightly different results from Datalog engines.
 The TC query returns correctly on all evaluated systems even though it uses bag semantics, even with cyclic input data.
@@
 If the query reaches a cycle in the data, the weight of the path will be infinitely increasing and the ever-changing cost column will prevent the set difference from removing already-discovered paths.
-The property responsible for this is constructor-freedom: the TC query constructs new values used to detect cycles even using bag semantics while the SSSP query constructs new values that lead to cyclic reinforcement and non-termination.
+The property responsible for both behaviors is constructor-freedom (P6): the TC query constructs new values used to detect cycles while the SSSP query constructs new values that lead to cyclic reinforcement and non-termination.
+The SSSP and TC queries exemplify why TyQL cannot strictly enforce all properties, even on a single RDBMS, as the same property can be responsible for nontermination in some queries but prevent nontermination in others.
 In summary, we evaluate the ability of TyQL to identify queries that will fail and find that it can be tuned to successfully reject all problematic queries.
+However, as there are queries that will run without problems that violate one or more of P1-P6, the strict safety guarantees of TyQL come at the cost of expressivity.
 As with many type systems, TyQL takes a conservative approach to correctness and can reject queries that may, if the data has certain properties, return successfully.
-To maximize usability and practicality, TyQL users always have the choice to tune how restrictive the fix operator is, potentially allowing the full RQB to compile.
+To maximize usability and practicality, TyQL users always have the choice to tune which combination of properties P1-P6 are relaxed, potentially allowing the full RQB to compile.
 8.3. Performance and State-of-the-Art
+In this section we evaluate TyQL with respect to performance.
+Application developers who wish to run fixed-point algorithms on data stored in a RDBMS have several options.
+The most immediate choice is to simply read data into memory and then execute their algorithm using the programming language constructs.
+Unfortunately, the Scala Collections API does not include a built-in fixed-point operator, so users must implement their own iterative control flow.
+The benefit of this approach is that it is fully customizable, yet it puts the burden onto the developer and the machine where the application is running.
+Alternatively, users may offload computation to the database.
+A natural way to do this would be to use language-integrated query to compose queries that are compiled to SQL and sent to the RDBMS.
+Yet if the query library does not support WITH RECURSIVE, then users will need to handle control flow at the application level and send only non-recursive queries at each iteration, or default to raw SQL strings.
 | Query | Size | It | TyQL (s) | Collections (s) | ScalaSQL (s) | SQL String (s) | vs. Collections | vs. ScalaSQL | vs. SQL String |
@@
 | 15MB | - | 1.324 | TO | 9.976 | 1.356 | ¿453.17X | 7.54X | 1.02X* |
-In this section we evaluate TyQL with respect to performance.
-Application developers who wish to run fixed-point algorithms on data stored in a RDBMS have several options.
-The most immediate choice is to simply read data into memory and then execute their algorithm using the programming language constructs.
-Unfortunately, the Scala Collections API does not include a built-in fixed-point operator, so users must implement their own iterative control flow.
-The benefit of this approach is that it is fully customizable, yet it puts the burden onto the developer and the machine where the application is running.
-Alternatively, users may offload computation to the database.
-A natural way to do this would be to use language-integrated query to compose queries that are compiled to SQL and sent to the database.
-Yet if the query library does not support WITH RECURSIVE, then users will need to handle control flow at the application level and send only non-recursive queries at each iteration, or default to raw SQL strings.
 8.3.1. Experimental Setup
@@
 The “ScalaSQL” column shows the execution time of the latest state-of-the-art language-integrated query library ScalaSQL, using non-recursive SQL queries (as recursion is not supported).
-This approach is representative of other language-integrated query libraries in Scala that support only non-recursive SQL.
+This approach is representative of other language-integrated query libraries in Scala since they support only non-recursive SQL.
 Because we use an embedded database system that runs within the same process as the application, avoiding the overhead of round-trips between database and application, this approach is equivalent to a PL/SQL approach.
@@
 The “TyQL” column shows the execution time of the query using TyQL.
-The final columns, “vs.
+The rightmost columns, “vs.
 Collections”, “vs.
@@
 The fixpoint implementation used in both the non-recursive SQL and the collections-only implementation is tail-recursive and based on the canonical example given in Scala by Example (Odersky, 2005) extended to use the same bottom-up Semi-Naive evaluation algorithm used internally in the database.
-The database used is in-memory DuckDB v1.1 with JDBC driver v.1.1.0 and queries that risked nontermination are run with set-semantics.
+The database used is in-memory DuckDB v1.1 with JDBC driver v.1.1.0 and queries that risked nontermination are run with set semantics.
 Each query is run on synthetic data of three different input relation sizes.
@@
 8.3.2. Experimental Results and Analysis
-| Slowdown vs. SQL String | Risks | ||||||
+The smallest of the three input data sets shows the range of data sizes for which the Collections API outperforms the other approaches due to avoiding the overhead of database connection and initialization.
+ScalaSQL has a higher overhead than TyQL or SQL strings, as the query initialization overhead happens at every iteration.
+The medium-sized dataset shows data input sizes where the ScalaSQL approach outperforms the Collections API due to several reasons: the RDBMS query optimizer can effectively select efficient query plans and join algorithms; as the data is not sent back to the application at each iteration, overheads due to boxing and unboxing of primitive types are avoided until iteration has concluded.
+For this input data size TyQL outperforms the other approaches due to avoiding multi-query overhead, storing and copying intermediate relations, and internal database optimizations.
+The memory usage of Collections is higher than either the TyQL or ScalaSQL because the data is not stored in the RDBMS, putting additional pressure on the JVM garbage collector.
+The largest data size shows the cases where the Collections API and non-recursive SQL approaches may run out of memory or time-out after 10-minutes.
+The graph algorithm queries (Ancestry, TC, SSSP) are run on larger datasets, while program analysis queries (APT, CBA, Data Flow, JPT, PTC) are run on smaller datasets to avoid all systems running out of memory.
+| Slowdown vs. SQL String | Risk Mitigation | ||||||
 | Max | Min | Avg | Data-type errors | Runtime DB error | Incorrect results | Nontermination | |
@@
 | TyQL | 1X | 1X | 1X | ✓ | ✓ | ✓ | ✓ |
-The smallest of the three input data sets shows the range of data sizes for which the Collections API outperforms the other approaches due to avoiding the overhead of database connection and initialization.
-ScalaSQL has a higher overhead than TyQL or SQL strings, as the query initialization overhead happens at every iteration.
-The medium-sized dataset shows data input sizes where the ScalaSQL approach outperforms the Collections API due to several reasons: the data management system query optimizer can effectively select efficient query plans and join algorithms; as the data is not sent back to the application at each iteration, overheads due to boxing and unboxing of primitive types are avoided until iteration has concluded.
-For this input data size TyQL outperforms the other approaches due to avoiding multi-query overhead, storing and copying intermediate relations, and internal database optimizations.
-The memory usage of Collections is higher than either the TyQL or ScalaSQL because the data is stored externally, and does not put any pressure on the JVM garbage collector.
-The largest data size shows the cases where the Collections API and non-recursive SQL approaches may run out of memory or time-out after 10-minutes.
-The graph algorithm queries (Ancestry, TC, SSSP) are run on larger datasets, while program analysis queries (APT, CBA, Data Flow, JPT, PTC) are run on smaller datasets to avoid all systems running out of memory.
 Statistics over the data presented in Table 4 are shown in Table 5.
@@
 TyQL shows no performance penalty compared to raw SQL strings, with a significant performance gain over Collections and ScalaSQL using non-recurive SQL.
-The options present a classic tradeoff between customizability and performance: raw SQL strings show the best performance but provide no safety guarantees, hand-written imperative programs provide expressibility and flexibility, language-integrated non-recursive queries put less burden on the developer and show better performance, and lastly TyQL puts the least burden on the developer and shows performance equivalent to raw SQL strings while providing safety guarantees.
+Table 5 illustrates the tradeoff between customizability, performance, and safety in the state-of-the-art for query execution: raw SQL strings show the best performance but provide no safety guarantees; hand-written imperative programs provide expressibility and flexibility, but only safety with respect to the programming language; language-integrated non-recursive queries put less burden on the developer and show better performance than imperative implementations and worse performance than raw SQL, while providing only non-recursive database safety guarantees, and lastly TyQL puts the least burden on the developer and shows performance equivalent to raw SQL strings while providing the strongest safety guarantees.
 9. Related Work
@@
 Novel extensions to SQL with cleaner recursion semantics have been proposed (Hirn and Grust, 2023) but are not widely implemented in commercial databases.
-Datalog is one of the most successful query languages with recursion capabilities, which has been used for fixpoint computations in program analysis (Scholz et al., 2016; Fan et al., 2019b; Jordan et al., 2016).
-As the core Datalog disallows non-monotone operations, various extensions of it have been proposed (Klopp et al., 2024b; Shaikhha et al., 2024; Klopp et al., 2024a; Wang et al., 2021).
+Datalog is one of the most successful query languages with recursion capabilities, which has been used for fixpoint computations in program analysis (Scholz et al., 2016; Fan et al., 2019; Jordan et al., 2016).
+As the core Datalog disallows non-monotone operations, various extensions of it have been proposed (Klopp et al., 2024; Shaikhha et al., 2024; Wang et al., 2021).
+(Starup et al., 2023) encodes dependency graphs in the type system with the goal of finding a safe stratification for Datalog with negation, while enforces stratification structurally, by how queries are composed in the host language.
 Flix (Madsen et al., 2016; Madsen and Lhoták, 2020) is a general-purpose language that supports Datalog queries as first-class citizens.
-Datafun (Arntzenius and Krishnaswami, 2016) is a functional Datalog variant that tracks monotonicity, constructor-freedom, and range-restriction via type-level constraints.
-Unlike TyQL, Flix and Datafun have their own runtime and execution engine based on Datalog semantics, while TyQL translates into a subset of safe SQL.
+Datafun (Arntzenius and Krishnaswami, 2016) is a functional Datalog variant that tracks range-restriction, monotonicity, and constructor-freedom via type-level constraints.
+Both Flix and Datafun have their own runtime and execution engine based on Datalog semantics.
+In contrast, is designed as a host-language-embedded DSL that generates recursive SQL targeting real-world databases with inconsistent semantics.
+The philosophy of is to be ”backend-polymorphic”, so instead of relying on a fixed operational semantics and built-in runtime, derives properties P1-P6 from recursive queries and rejects queries that risk the unwanted database behaviors B1-B3.
 10. Conclusion
@@
 TyQL provides a clean abstraction for recursive queries while ensuring correctness and safety at compile-time that is specialized to the database.
-We formalized the constraints with , which prevents unwanted behavior across databases, including errors, incorrect results, and nontermination.
+We formalized the constraints with , which prevents database runtime errors, incorrect results, and nontermination.